### v1

# LimiX: Unleashing Structured-Data Modeling Capability for Generalist Intelligence

Abstract
We argue that progress toward general intelligence requires complementary foundation models grounded in language, the physical world, and structured data. This report presents LimiX, the first installment of our large structured-data models (LDMs). LimiX treats structured data as a joint distribution over variables and missingness, thus capable of addressing a wide range of tabular tasks through query-based conditional prediction via a single model. LimiX is pretrained using masked joint-distribution modeling with an episodic, context-conditional objective, where the model predicts for query subsets conditioned on dataset-specific contexts, supporting rapid, training-free adaptation at inference. We evaluate LimiX across 10 large structured-data benchmarks with broad regimes of sample size, feature dimensionality, class number, categorical–to-numerical feature ratio, missingness and sample-to-feature ratios. With a single model and a unified interface, LimiX consistently surpasses strong baselines including gradient-boosting trees, deep tabular networks, recent tabular foundation models, and automated ensembles, as shown in Figure 1 and Figure 2. The superiority holds across a wide range of tasks, such as classification, regression, missing value imputation, and data generation, often by substantial margins, while avoiding task-specific architectures or bespoke training per task. All LimiX models are publicly accessible under Apache 2.0.
Figure 1: Performance comparison on the averaged reciprocal of the ranks, where the rank is that of the corresponding model on ROC AUC. Higher values indicate stronger average ranking performance across all the classification benchmarks.
Figure 2: Performance comparison on the averaged reciprocal of the ranks, where the rank is that of the corresponding model on R2 across all the regression benchmarks.
1 Introduction
We posit that progress toward general intelligence is best organized around three complementary spaces: language, physical-world, and structured data, each anchored to a distinct data modality and set of inductive biases. In the language space, large language models (LLMs) provide a universal interface for natural and programming languages and have rapidly advanced instruction following, tool use, and explicit reasoning over token sequences (Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023; Bai et al., 2023). In the physical-world space, recent foundation models ground knowledge in space perception and embodied reasoning. They emphasize spatial intelligence through structured scene understanding, controllable scene generation, and neural radiance field reconstruction (Mildenhall et al., 2020; Ke et al., 2025; Xiang et al., 2025; Li et al., 2024a; Yi et al., 2024). These models also include self-supervised video world models such as V-JEPA, which learn predictive abstractions from large-scale videos and support downstream planning and control (Bardes et al., 2024; Assran et al., 2025).
On the other hand, structured data serves as the foundational bedrock for evidence-based decision-making across a multitude of critical domains, including finance, healthcare, logistics, and public policy (Fuster et al., 2022; Johnson et al., 2016; Yu et al., 2021; Krafft et al., 2020). The structural consistency and inherent order of structured data provide a robust framework for quantitative analysis and reliable operations (Ramakrishnan et al., 2003; Silberschatz et al., 2011; Stonebraker & Çetintemel, 2018), enabling precise prediction, automated reasoning, and rigorous causal inference (Pearl, 2009; Hernán & Robins, 2010; Little & Rubin, 2019). While the emergence of unstructured data has captured considerable attention, the analytical power and operational reliability of structured data remain unparalleled for a vast array of real-world applications (Fang et al., 2024; Van Breugel & Van Der Schaar, 2024). Consequently, advancements in structured-data prediction, analysis, and reasoning are not merely an academic pursuit but a critical enabler for efficiency, innovation, and accuracy in modern data-driven systems. Structured data is not subsumed by language models or embodied intelligence. Converting tables into free text discards metric geometry, physical units, and patterns of missingness that are central to reliable prediction, while models focusing on perception and control in three-dimensional physical environments do not capture discrete interventions, business rules, or causal heterogeneity across environments. Empirical surveys also document current limitations of language models on tabular prediction without bespoke adaptation (Fang et al., 2024; Sui et al., 2024).
Traditionally, practitioners deploy pipelines of specialized models with gradient-boosting trees and automated ensembles such as XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), CatBoost (Dorogush et al., 2018), and AutoGluon (Erickson et al., 2020)—that are trained separately for each dataset and task. These systems excel at supervised prediction but require full retraining on every new dataset, prolonging deployment and preventing reuse of knowledge across domains. Recent deep approaches for tables have improved accuracy on mixed-type data (Huang et al., 2020; Yoon et al., 2020; Gorishniy et al., 2021; Somepalli et al., 2022; Bahri et al., 2022). However, they are still typically trained per dataset and do not provide a single model that transfers across objectives and constraints.
These limitations motivate the pursuit of a foundation model for structured data, i.e. models trained on large families of datasets to perform in-context learning (ICL) without per-task fine-tuning. Notably, TabPFN (Hollmann et al., 2022) and its successor TabPFN-v2 (Hollmann et al., 2025) demonstrate state-of-the-art performance and speed on small-to-medium-scale tables via prior-data fitting over diverse generative processes, and community efforts increasingly argue for tabular foundation models as a distinct paradigm (Van Breugel & Van Der Schaar, 2024). Concurrent works such as TabDPT (Ma et al., 2024) and TabICL (Qu et al., 2025) explore scaling these ideas to real and larger datasets, narrowing the gap between tabular foundation models and classical methods under the scenarios of larger sample sizes.
Despite the progress, these are still in the early stages of foundation model development and remain limited in generality and performance. Most models are developed and evaluated primarily for supervised prediction (classification or regression), and typically require per-task models, adapters, or external pipelines to address other objectives. In practice, one still needs to assemble separate components for classification, regression, missing value imputation, data generation, and sample selection for interpretability, with different training losses and hyperparameters, so the resulting system is not a single reusable learner that delivers all of these capabilities end-to-end while maintaining reliable performance. This gap motivates a large structured-data model (LDM) that treats structured data as a joint distribution over variables and missingness, enabling multiple tasks to be posed as queries to one model.
In this work, we introduce LimiX, the first installment of our LDM series. LimiX aims to push generality further: a single model capable of classification, regression, missing value imputation, data generation, and sample selection for interpretability under one training and inference recipe, shifting the paradigm from bespoke pipelines to unified, foundation-style tabular learning. LimiX adopts a lightweight, scalable architecture that represents structured data as a set of sample–feature embeddings and learns dependencies across two dimensions: across features (columns) and across samples (rows). To make the attention module explicitly column-aware without inflating parameters, we introduce a low-rank discriminative feature encoding that encodes feature identities. Pretraining of LimiX follows a masked joint-distribution objective and an episodic, context-conditional formulation: For each dataset, an in-context subset establishes dataset-specific priors, and the model is trained to predict masked entries in a disjoint query subset, enabling per-dataset adaptation without fine-tuning at inference. The pretraining corpus consists of data synthesized from hierarchical structural causal models (SCMs). Within the synthesis pipeline, we employ graph-aware sampling to obey the causal structure and solvability-aware sampling to accommodate the data quantity of various downstream tasks, improving coverage of local patterns and generalization. At inference, attention-guided retrieval provides an efficient, optional ensemble and fine-tuning mechanism. LimiX retrieves informative samples and features using its own attention scores, aggregates predictions across a handful of lightweight pipelines, and delivers calibrated outputs for various downstream tasks, all through a unified conditional-inference interface and without task-specific architectures or bespoke per-dataset training.
We evaluate LimiX on 10 large structured-data benchmarks with broad regimes of sample size, feature dimensionality, number of classes, categorical–to-numerical feature ratios, missing values, and sample-to-feature ratios. Results show that LimiX demonstrates the strongest predictive performance on all the benchmarks. With a single model and a unified inference interface, LimiX surpasses competitive baselines including gradient-boosting trees, deep tabular networks, recent tabular foundation models, and automated ensemble methods. Across classification, regression, missing value imputation, data generation, and out-of-distribution prediction, LimiX delivers consistent gains, often by large margins, while avoiding task-specific model architectures, customized ensembles, or per-dataset training. On most benchmarks, such as OpenML-CC18 (Bischl et al., 2017), TabArena (Erickson et al., 2025), TALENT-REG (Liu et al., 2024), LimiX is the only model that consistently outperforms AutoGluon, which is considered an outstanding baseline across various tabular-data tasks.
2 Architecture
We consider a dataset $\mathcal{D}=\{({\mathbf{x}}^{R},{\mathbf{y}}^{R})\}$ of $d$ features and an outcome variable, where ${\mathbf{x}}^{R}=\{{\mathbf{x}}^{R}_{i}\}_{i=1}^{m}$, ${\mathbf{y}}^{R}=\{y^{R}_{i}\}_{i=1}^{m}$ and the superscript $R$ indicates the raw input. Here, ${\mathbf{x}}^{R}_{i}\in\mathbb{R}^{d}$ and $y^{R}_{i}\in\mathbb{R}$ correspond to the $i^{th}$ sample, while ${\mathbf{x}}^{R}\in\mathbb{R}^{m\times d}$ and $\mathbf{y}^{R}\in\mathbb{R}^{m}$ correspond to the 2D tabular data. For in-context samples and test samples, we use subscripts to distinguish between them. For example, ${\mathbf{x}}^{R}_{ct}\in\mathbb{R}^{m_{ct}\times d}$ denotes the features of in-context samples, while ${\mathbf{x}}^{R}_{te}\in\mathbb{R}^{m_{te}\times d}$ denotes those of the test samples.
2.1 Embedding of Tabular Data
To ensure compatibility with modern architectures such as Transformers (Vaswani et al., 2017), each cell of the 2D tabular input $x^{R}_{i,j}\in\mathbb{R}$ is first projected into a latent embedding space $\mathbb{R}^{p}$. Specifically, ${\mathbf{x}}^{R}\in\mathbb{R}^{m\times d}$ is transformed into ${\mathbf{x}}\in\mathbb{R}^{m\times d\times p}$ and $\mathbf{y}^{R}\in\mathbb{R}^{m}$ is transformed into ${\mathbf{y}}\in\mathbb{R}^{m\times p}$. All subsequent attention operations are then conducted within this latent embedding space. Such a high-dimensional embedding space could strengthen the expressivity of the model. Concretely, we employ a two-layer MLP with LayerNorm (Ba et al., 2016) as the embedding module, i.e. ${\mathbf{x}}_{i,j}=\sigma(LN(\sigma(LN({\mathbf{x}}^{R}_{i,j}{\mathbf{W}}^{(1)}+{\mathbf{b}}^{(1)})){\mathbf{W}}^{(2)}+{\mathbf{b}}^{(2)}))$, where LN is LayerNorm and $\sigma$ is the GELU activation function (Hendrycks & Gimpel, 2016). Separate embedding modules are used for ${\mathbf{x}}^{R}$ and ${\mathbf{y}}^{R}$.
Figure 3: The overall model structure of LimiX.
2.2 Discriminative Feature Encoding
The attention score between feature $j$ and $j^{\prime}$ for sample $i$ is $\frac{1}{\sqrt{p}}\left\langle{\mathbf{x}}_{i,j}{\mathbf{W}}_{\mathrm{Q}},{\mathbf{x}}_{i,j^{\prime}}{\mathbf{W}}_{\mathrm{K}}\right\rangle$. This score depends solely on the interactions between the embeddings of cell values and imposes no explicit prior on feature (column) identity. As a result, the model cannot reliably infer the column from which a cell value originates. Inspired by Hollmann et al. (2025), we introduce a learnable low-rank column identifier termed discriminative feature encoding (DFE) whose design philosophy includes two objectives: (i) Encodings of different features should be well separated to ensure discriminability; (ii) Encodings in the embedding space admit a low effective rank so that the model can express column identities compactly and share statistical strength across features.
For implementation, let $s$ denote the rank of DFE, and $s\ll p$ compared with the dimension of the embedding space $p$. We initialize a matrix ${\mathbf{u}}\in\mathbb{R}^{d\times s}$ whose $j^{th}$ row vector ${\mathbf{u}}_{j}\in\mathbb{R}^{s}$ is a low-dimensional code of the column identifier corresponding to feature (column) $j$. Rows are initialized to be approximately orthogonal and normalized. Then a linear transformation ${\mathbf{E}}\in\mathbb{R}^{s\times p}$ lifts the code from the low-dimensional code space to the high-dimensional embedding space, i.e. ${\mathbf{e}}_{j}=\,{\mathbf{u}}_{j}{\mathbf{E}}\in\mathbb{R}^{p}$, serving as the DFE for feature $j$. Finally, the embedding of cell $(i,j)$ is augmented additively as $\tilde{{\mathbf{x}}}_{i,j}={\mathbf{x}}_{i,j}+{\mathbf{e}}_{j}$, which is analogous to absolute positional encodings but is applied along the feature axis. Empirically, we set $s=p/4$ by default.
2.3 Model Structure
As shown in Figure 3, LimiX comprises 12 transformer blocks. Each block performs axis-wise self-attention along the feature axis twice and along the sample axis once, and each attention sublayer is followed by a position-wise feed-forward network (FFN). We adopt this asymmetric design with two feature-level passes and one sample-level pass because ablations show that using equal numbers underrepresent feature interactions, whereas doubling attention along the feature axis improves modeling capacity on heterogeneous schemas with minimal overhead. All sublayers use LayerNorm (Ba et al., 2016) in a pre-normalization configuration to stabilize optimization and support scaling in depth.
3 Pretraining
3.1 Context-Conditional Masked Modeling for Joint Distribution Learning
We pretrain LimiX by randomly masking cells in each row and training the model to recover the hidden entries from the visible context. Exposing the model to various mask patterns forces it to master a wide spectrum of conditional dependencies among variables. When these conditionals are learned properly, they effectively define a single joint model of the data. This joint model can then be queried for diverse tasks with one mechanism: Treat a chosen column as the target to perform regression or classification, fill in missing values by predicting the masked cells from the observed ones, and generate new samples by iteratively masking and refilling subsets of features.
To better align pretraining with inference, we adopt a Context-Conditional Masked Modeling (CCMM) objective. For each dataset, an episode splits rows into a context subset and a query subset. The model encodes the context and conditions predictions for the query rows on this context through attention or feature modulation, learning to answer queries of the form “predict masked entries in the query given the observed query features and the context”. This episodic formulation enables rapid and label-free adaptation to new datasets at inference time: A handful of context rows establish dataset-specific priors such as category frequencies, marginal scales, and cross-feature couplings, while a single parametric model serves all schemas and tasks.
Unlike BERT-style masked modeling (Devlin et al., 2019), where adaptation is implicit in parameters accumulated during pretraining, CCMM considers context as a non-parametric memory that the model can consult during inference. This yields per-dataset calibration, better handling of rare categories, and improved robustness to distribution shift without gradient updates, aligning with principles of in-context learning and meta-learning for mixed-type tabular data.
We also find that CCMM improves modeling of the underlying dependency structure compared with recent tabular foundation models (Hollmann et al., 2025; Qu et al., 2025). By demanding consistency across numerous conditional predictions of all features rather than optimizing the loss of a single conditional prediction of a prefixed feature, the model is compelled to capture stable variable–variable relations instead of brittle decision boundaries. As the coverage of masks becomes richer, these cross-conditional constraints get tightened, yielding more reliable recovery of the joint distribution of all the features and more stable estimates in downstream usages. Please refer to Section 6 for details.
3.2 Mask Pattern Design
We employ a heterogeneous schedule that interleaves cell-wise, column-wise, and block masks, enabling the model to practise recovering both isolated entries and coordinated subsets of variables. Cell-wise masks refine local conditional predictions; column-wise masks force the model to treat an entire feature as missing and infer it from the remaining attributes; block masks target higher-order dependencies across semantically related groups (e.g., demographics with outcomes, laboratory panels with diagnoses). Masking rates are stratified by variable type, prevalence, and dispersion to avoid overfitting to common categories and to limit the influence of high-variance continuous features, and we exclude degenerate patterns that remove nearly all informative context. This diversified schedule provides broad coverage of conditional relationships and prevents the model from specializing to a narrow reconstruction regime, yielding a more faithful approximation of the joint distribution. In practice, we sample the mask ratio in [0.1, 0.4] for LimiX.
3.3 Mask Embedding
To model masked cells, we introduce learnable mask embeddings that explicitly mark positions to be predicted. For each masked entry, its embedding is replaced by a trainable mask vector that is combined with the column embedding, so the encoder can condition on what is missing as well as where. We align the training masks introduced by the objective with naturally missing/structurally unavailable values observed in real data by using the shared mask embedding and calibrating the masking schedule to match empirical missingness patterns, thereby minimizing distribution shift between synthetic and real scenarios.
The mask embeddings flow through the same attention blocks as observed cells, enabling the model to request information from relevant columns and to produce calibrated distributional predictions at the output head. To reduce the mismatch between pretraining and fine-tuning, we condition the network on the corruption level of each dataset using a mask density feature, defined as the proportion of cells that are masked in the dataset. This scalar is encoded via a lightweight statistics token whose embedding is conditioned on the mask density with a small MLP module. The mask ratio embedding is included alongside the data tokens and participates in self-attention. Providing this cue regularizes the model across a range of masking rates and reduces pretraining to inference mismatch, which improves calibration when the inference pattern, such as masking a single target column, differs from the heavier masks used during pretraining.
4 Pretraining Data Generation
The performance of foundation models largely depends on the diversity and quality of pretraining data. To obtain a well-generalized foundation model for tabular data, we generate synthetic datasets using Directed Acyclic Graphs (DAGs), enabling the creation of datasets with diverse characteristics. The data generation process consists of three stages: DAG generation, data sampling, and task adaptation. First, a DAG is constructed to represent complex causal dependencies among variables. Then, a subset of these variables is sampled to define a specific problem, allowing LimiX to develop causal reasoning capabilities. Finally, the sampled data is processed to align with various downstream tasks. Compared with the data generation methods used in current foundation models  (Hollmann et al., 2025; Qu et al., 2025), we adopt a hierarchical generation paradigm that establishes causal dependencies in a more controllable and interpretable manner, while the sampling strategies further enhance the ability to generate datasets with diverse solvability and characteristics.
4.1 DAG Generation based on Hierarchical SCMs
The theoretical foundation of data generation lies in structural causal models (SCMs). As illustrated in Figure 4, within a DAG, initial data for each root node is independently sampled from an assigned distribution, with the distributions themselves chosen from a collection parameterized by randomized hyperparameters. Beginning at the root nodes, the initial data propagates through the DAG along its directed edges, wherein each encountered node represents a distinct local causal structure (LCS). As previously mentioned, the diversity of synthetic data is crucial for effective pretraining. To ensure that DAG generation is both diverse and well-structured, we adopt a hierarchical generation scheme rather than generating DAGs directly, thereby allowing for more fine-grained controls over the generation process. Within the LCS, the data input first propagates to the connected parent nodes, and the data of the child node can be obtained through $X_{i}=f(\{g_{i,k}(X_{k})_{k\in{\rm PA}(X_{i})},\varepsilon_{i})\}$ where $g_{i,k}(\cdot)$ represents the edge function from $X_{k}$ to $X_{i}$, ${\rm PA}(X_{i})$ denotes the set of parent nodes of $X_{i}$, $f(\cdot)$ is a parameterized aggregation function, and $\varepsilon_{i}$ is an observational noise. This framework allows us to capture diverse local causal dependencies. For example, if an LCS contains only one parent node, that node is the direct cause of the child, making the dependency clear and straightforward. However, when multiple parent nodes are included in the LCS, the causal relationship becomes more complex due to the presence of potential confounding variables within the network.
Figure 4: An example of the generated DAG, where $g_{i,j}$ is an edge function defining the relationship between a parent node $X_{i}$ and its child node $X_{j}$ in the local causal structures.
In addition to the structural properties of LCSs, edge functions and the aggregation function also play crucial roles in shaping the dependencies. In this work, we use three types of edge functions:
  • Multilayer perceptrons (MLPs): To determine the architecture of these neural networks, we uniformly sample properties such as the number of linear layers and their associated activation functions. When multiple layers are included, the MLP introduces complex nonlinear dependencies along the edge, whereas with only a single layer, it degenerates into a simple transformation. For weight initialization, we randomly choose from Xavier initialization (Glorot & Bengio, 2010) and He initialization (He et al., 2015). The activation functions are uniformly sampled from identity mapping, hyperbolic tangent, sigmoid, logarithm, absolute value, sine, squaring, modulo operation, heaviside step function, and GELU.
  • Convolutional layers: For node-level tabular data, the convolution operation is applied along the sample dimension, serving as a local information mixer. The choices of weight initialization and activation functions follow the same procedure as those employed for MLPs.
  • Decision trees: These are employed to introduce rule-based mappings and can take the form of either classification or regression trees. Unlike some approaches that fit decision trees on random data, we argue that model fitting is neither necessary nor computationally efficient for the purpose of introducing rule-based dependencies. In this work, decision trees are fixed once constructed, with their hyperparameters sampled independently for each edge.
Once the child node receives the mapped values, they are aggregated using an aggregation function selected from one of three options: simple average, weighted average, or MLP-based aggregation. With respect to observational noise introduced at each edge, rather than adding noise of a fixed magnitude, we generate noise whose magnitude is scaled according to the distribution of each feature. To construct the DAG from all LCSs, we begin with a single-LCS DAG. Each time a new LCS is incorporated, it may connect to one or multiple existing LCSs in the DAG, whereby the child node of the new LCS is linked to the parent nodes of the target LCSs. This process is repeated until no LCSs remain to be added. In the resulting structure, all nodes with an in-degree of zero serve as the root nodes.
4.2 Data Sampling
Determining how to sample a subset of the generated data as the training set is also a critical challenge. Although random sampling is possible, the resulting training data often deviates significantly from being truly representative or good for training models. Thus it is necessary to devise an efficient strategy to sample high-quality training data for pretraining. We employ two sampling strategies: graph-aware sampling and solvability-aware sampling. Graph-aware sampling can be regarded as a more advanced variant of random sampling, where the key difference lies in its consideration of the graphical distribution of the sampled training data, which constrains the sampling space and renders it significantly smaller than in the case of random sampling. In contrast to graph-aware sampling, solvability-aware sampling aims to provide training data with varying degrees of solvability, thereby encouraging the model to achieve better generalization. From this point, we divide the subsampled problems into three classes: high-solvability, moderate-solvability and low-solvability problems. We ensure that the sampling ratios of these classes follow a categorical distribution, each of whose parameters is sampled from a distinct Gaussian distribution. In practice, we alternate between the two sampling strategies according to a predefined probability.
4.3 Task Adaptation
During the data generation process, the sampled data may be intended for either classification or regression tasks, and the major difference is the processing of the target variable $y$, which is initially sampled as a continuous variable. For classification tasks, it is subsequently discretized into categorical variables. For regression tasks, the procedure differs slightly. If the sampled $y$ is already a discrete variable, a regression task will not be constructed; In cases where $y$ is continuous but clusters closely around a limited set of distinct values, an in-order transformation can be applied to achieve a more uniform distribution across the magnitude scale.
5 Retrieval-based Ensemble
We adopt an inference-time and retrieval-based ensemble strategy that leverages LimiX's learned attention scores to upweight and select representative in-context samples and features without any additional training, so that we can further improve the performance of LimiX.
In terms of ensemble, we run multiple inference pipelines per dataset and aggregate the results. In each pipeline, we (i) randomly permute the feature columns and reorder the labels for categorical features or outcomes, and (ii) augment a subset of features with simple, schema-preserving transformations like quantile normalization, log-normal transformation, and high-energy SVD components. For classification, the number of inference pipelines is set to 4. For regression, it is set to 8.
In terms of retrieval, we perform two forward passes of LimiX for each pipeline. The first pass is performed based on all in-context samples and is employed for retrieval. The second pass is performed based on the customized in-context samples retrieved in the first pass.
The procedure of the first pass is as follows. First, last-layer feature-level attention provides importance scores over features, which can be used as feature weights for subsequent sample selection. Concretely, for each test sample, we calculate ${\mathbf{a}}_{f}\in\mathbb{R}^{d+1}$, which is the feature-level attention score between the outcome $y_{te}$ and the concatenation of $F$ features and outcome $({\mathbf{x}}_{te},y_{te})$. Then, the module of last-layer sample-level cross-attention between test samples and in-context samples induces importance scores over in-context samples. Concretely, we calculate ${\mathbf{a}}_{s}\in\mathbb{R}^{m_{ct}\times(d+1)}$, which is the sample-level cross-attention scores between the test sample $({\mathbf{x}}_{te},y_{te})$ and in-context samples $({\mathbf{x}}_{ct},{\mathbf{y}}_{ct})$. Finally, for a given test sample, we retrieve in-context samples with highest reweighted attention scores as customized in-context samples for this test sample, and perform a forward propagation again for prediction. Concretely, we calculate ${\mathbf{a}}_{sf}={\mathbf{a}}_{s}{\mathbf{a}}_{f}\in\mathbb{R}^{m_{ct}}$, a weighted average of ${\mathbf{a}}_{s}$ along the feature dimension with ${\mathbf{a}}_{f}$ as the feature weights.
(a) Visualization of synthetic samples. Each sector represents a category of in-context samples. The yellow pentagram represents the test sample.
(b) Top 10% of in-context samples sorted by sample-level attention scores.
Figure 5: Toy example of sample-level attention. In-context samples that share the same category as the query sample are assigned higher scores through the attention module of LimiX.
(a) Causal DAG of synthetic samples.
(b) Heatmap of feature-level attention scores of $Y$.
Figure 6: Toy example of feature-level attention for the outcome variable. Direct causes of the outcome and the outcome itself are assigned almost all the attention weights in LimiX.
Toy examples.
We present the effect of bi-level attention-based retrieval via toy examples.
  • Sample-level retrieval. We generate 2D synthetic data of 10 classes, where each class of data points occupies a sector of a circle in Figure 5(a). Since attention depicts sample similarity in the latent embedding space and it also takes the dependency between input features and the category label into account, it is capable of capturing more complex dependencies than naively using Euclidean distance in the original 2D space of features. From Figure 5(b), we can see that the attention module predominantly assigns large weights to in-context instances sharing the same category label as the query instance. This indicates that the model leverages class-consistent contextual information from in-context samples to assist its prediction.
  • Feature-level retrieval. We generate synthetic data via the SCM in Figure 6(a), where each green node denotes an observed feature and the yellow node denotes the outcome. Each edge is a two-layer MLP using ReLU as the activation function added with Gaussian noise. From Figure 6(b), we can see that the attention module assigns most weights to a subset of features, which is exactly the set of direct causes of the outcome in the SCM. This suggests that feature-level attention could help to focus on causal features and reduce reliance on spurious correlations.
6 Theoretical Analysis on Context-Conditional Masked Modeling
We begin by introducing the mathematical formulation underlying our theoretical analysis.
Notations.
Let $\Omega$ denote the space for each feature in the table. We represent the $m$ in-context samples, each with $d$ features, by the random matrix $\mathbf{X}^{\text{ct}}\in\Omega^{m\times d}$. We use the random vector $\mathbf{X}^{\text{te}}=(X^{\text{te}}_{1},\dots,X^{\text{te}}_{d})\in\Omega^{d}$ to denote the test sample111 For simplicity, we assume (i) all features share the same space $\Omega$, though the results generalize to distinct spaces; (ii) we use $m$ instead of $m_{ct}$ in Section 2; and (iii) there is only one test sample, i.e., $m_{te}=1$ in Section 2.. Meanwhile, we do not explicitly introduce a target label $y$ in this section for simplicity; instead, we treat it as a particular dimension of both the in-context and test samples. Consequently, one dimension of $\mathbf{X}^{\text{te}}$ is unknown to the model during the test phase.
For any $\pi\subseteq[d]$, let $\mathbf{X}^{\text{te}}_{\pi}=(X^{\text{te}}_{j})_{j\in\pi}$ and $\mathbf{X}^{\text{te}}_{-\pi}=(X^{\text{te}}_{j})_{j\notin\pi}$ denote the subvectors on $\pi$ and its complement, respectively. In particular, for $\pi=\{j\}$ we use $\mathbf{X}^{\text{te}}_{-j}$ as shorthand for $\mathbf{X}^{\text{te}}_{-\{j\}}$.
Training and test procedures.
We are given $n$ i.i.d. samples $\{(\mathbf{x}^{\text{ct},(i)},\mathbf{x}^{\text{te},(i)})\}_{i=1}^{n}$ drawn from $p(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})$, given in Section 4. The model is trained using CCMM, a masked pre-training method (Section 3). Let $\Pi\subseteq 2^{[d]}$ be a set of masks and $\mathrm{Unif}(\Pi)$ the uniform distribution over $\Pi$. In practice, we take $\Pi=\{\pi\subseteq[d]:|\pi|\in[0.1d,0.4d]\}$, while for theory we simplify to masks of fixed size $k$:
$\Pi_{k}:=\{\pi\subseteq[d]:|\pi|=k\}.$ (1)
It is easy to verify that our theoretical insights extend naturally to settings with variable mask sizes.
For each sample $(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})$, we draw $\pi\sim\mathrm{Unif}(\Pi_{k})$ and train $q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})$, where $\theta\in\Theta$ and $\Theta$ denotes the parameter space of the model, to reconstruct the masked features. Let $\pi_{i}$ be the mask for $(\mathbf{x}^{\text{ct},(i)},\mathbf{x}^{\text{te},(i)})$. The empirical loss $\hat{L}_{k}(\theta)$ and estimator $\hat{\theta}_{k,n}$ are
$\hat{L}_{k}(\theta)=\frac{1}{n}\sum_{i=1}^{n}-\log q_{\theta}\left(\mathbf{x}^{\text{te},(i)}_{\pi_{i}}|\mathbf{x}^{\text{te},(i)}_{-\pi_{i}},\mathbf{x}^{\text{ct},(i)}\right),\qquad\hat{\theta}_{k,n}=\operatorname*{arg\,min}_{\theta\in\Theta}\hat{L}_{k}(\theta).$ (2)
As $n\to\infty$, this converges to the population-level loss $L_{k}(\theta)$ and the corresponding solution $\theta^{*}_{k}$:
$L_{k}(\theta)=\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\,\pi\sim\mathrm{Unif}(\Pi_{k})}\left[-\log q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right],\qquad\theta^{*}_{k}=\operatorname*{arg\,min}_{\theta\in\Theta}L_{k}(\theta).$ (3)
At test time, all features in $\mathbf{X}^{\text{ct}}$ are observed, while one feature $X^{\text{te}}_{j}$ of $\mathbf{X}^{\text{te}}$ is missing and must be inferred. The goal is to output $p(X^{\text{te}}_{j}\mid\mathbf{X}^{\text{te}}_{-j},\mathbf{X}^{\text{ct}})$.
6.1 Modeling the Joint Conditional with Random Masks
We now explain the necessity of randomly sampling masks from $\mathrm{Unif}(\Pi_{k})$. Our result is based on the following proposition.
Proposition 6.1 (Informal; See Proposition B.1).
Under mild assumptions, for any $k\in[d]$, there is a one-to-one correspondence between the distribution $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ and the family of conditionals $\{p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}):\forall\pi\in\Pi_{k}\}$.
At test time, the target variable $y$ may correspond to any feature $X^{\text{te}}_{j}$ of the test sample $\mathbf{X}^{\text{te}}$, so the model must be able to estimate $p(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-j},\mathbf{X}^{\text{ct}})$ for every $j\in[d]$. As shown in Proposition 6.1, this requirement is equivalent to learning the joint conditional distribution $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$. Hence, a model can achieve strong predictive performance at test time if and only if it learns $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$. Moreover, if each $p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})$ can be well approximated by $q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})$ for all $\pi\in\Pi_{k}$, then the model can recover the joint conditional distribution and thereby generalize effectively. By contrast, Example B.1 shows that knowledge of $p(X^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})$ for only a subset of $\pi\in\Pi_{k}$ may be insufficient to recover $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$, preventing the model from generalizing effectively at test time.
Moreover, the target distribution $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ is closely connected to the underlying structural causal models (SCMs). Let $S$ denote a random variable corresponding to a structural causal model (SCM). By the data-generating process, we have $\mathbf{X}^{\text{ct}}\perp\mathbf{X}^{\text{te}}\mid S$, which yields
$p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})=\int_{S}p(\mathbf{X}^{\text{te}},S|\mathbf{X}^{\text{ct}})\,\mathrm{d}S=\int_{S}p(S|\mathbf{X}^{\text{ct}})\,p(\mathbf{X}^{\text{te}}|S,\mathbf{X}^{\text{ct}})\,\mathrm{d}S=\int_{S}p(S|\mathbf{X}^{\text{ct}})\,p(\mathbf{X}^{\text{te}}|S)\,\mathrm{d}S.$ (4)
Intuitively, Equation 4 suggests that one efficient way to learn $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ is through two components: (i) $p(S|\mathbf{X}^{\text{ct}})$, the posterior distribution over SCMs given the context, and (ii) $p(\mathbf{X}^{\text{te}}|S)$, the likelihood of the test sample under a given SCM.
6.2 The Choice of Mask Number
In this subsection, we demonstrate that the choice $k$ in Equation 1 has great impact on models' performances and we choose $k>1$ due to both sample efficiency and generalization considerations. Our result is an extension from the analysis by Li et al. (2024c) on the properties of masked sequence prediction.
Based on Proposition 6.1, we will henceforth use $q_{\theta}(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ to denote the distributions induced by the learned family of conditional probabilities $\{q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}):\pi\in\Pi_{k}\}$.
Sample efficiency.
Theorem below shows that larger $k$ yields lower estimation uncertainty.
Theorem 6.2 (Informal; see Theorem B.2).
Suppose there exists $\theta^{*}\in\Theta$ such that $q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})=p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})$ for all $\mathbf{X}^{\text{ct}}\in\Omega^{m\times d}$, $\mathbf{X}^{\text{te}}\in\Omega^{d}$, and $\pi\subseteq[d]$, and that the minimizer of $L_{k}(\theta)$ is unique for every $k$. Then, under mild regularity conditions, as $n\to\infty$,
$\sqrt{n}\,(\hat{\theta}_{k,n}-\theta^{*})\xrightarrow{d}{\mathcal{N}}(0,\Gamma_{k}),$
where $\Gamma_{k}$ does not depend on $n$ and satisfies $\Gamma_{k+1}\preceq\Gamma_{k}$.
The assumption on $\theta^{*}$ states that the optimal solution of the model can approximate any conditional distribution, which is reasonable given the expressive power of transformer-based architectures. This theorem further shows that for sufficiently large and fixed $n$, if $k_{1}>k_{2}$ then $\Gamma_{k_{1}}\preceq\Gamma_{k_{2}}$, implying that $\hat{\theta}_{k_{1},n}$ has lower estimation uncertainty than $\hat{\theta}_{k_{2},n}$. Equivalently, when $k$ is larger, fewer samples are required to achieve the same uncertainty, leading to greater sample efficiency.
Generalization for joint distribution learning.
Masked pretraining empowers the model to infer the joint distribution of data from the context of training samples. We further show that an increase in the density of random masks enables a more accurate reconstruction of the joint distribution.
Theorem 6.3 (Informal; See Theorem B.5).
Under regularity conditions on $q_{\theta}$, with high probability, for any $\theta\in\Theta$, the expected total variation between $q_{\theta}(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ and the true conditional distribution $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ is at most:
$\sqrt{\frac{1}{2}C_{k}(q_{\theta})\big{(}\hat{L}_{k}(\theta)+\text{complexity terms}\big{)}+{\mathcal{O}}(n^{-1/2})},$
where constants $C_{k}(q_{\theta})$ depend only on $q_{\theta}$ and $k$. Furthermore, $C_{k+1}(q_{\theta})\leq C_{k}(q_{\theta})$ for any $\theta\in\Theta$.
The theorem establishes an upper bound for the generalization error of the estimated conditional joint distribution $q_{\theta}(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ compared to the true joint distribution $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$, and shows that the upper bound decreases monotonically with respect to the number of masked cells.
7 Evaluation
7.1 Classification
Benchmarks.
For the quantitative evaluation of classification performance, we utilize multiple benchmarks, including TALENT-CLS (Liu et al., 2024), OpenML-CC18 (Bischl et al., 2017), PFN-CLS (Hollmann et al., 2025), TabZilla (McElfresh et al., 2023), and TabArena (Erickson et al., 2025). Among these benchmarks, the datasets containing more than 50,000 training samples (The number of testing samples is not constrained), 10,000 features, or 10 target categories were excluded. This selection process resulted in a final collection of 179 datasets from TALENT-CLS, 62 from OpenML-CC18, 29 from PFN-CLS, 27 from TabZilla, and 33 from TabArena.
Furthermore, we introduce Balanced Comprehensive Challenging Omni-domain (BCCO) Benchmark, comprising BCCO-CLS and BCCO-REG, for the evaluation of LimiX and baseline models. The BCCO benchmark is constructed from extensive open-source structured-data corpora, meticulously deduplicated and cleaned. It presents a significant challenge due to several intrinsic characteristics: the distribution of dataset attributes (such as the ratio of categorical features), and the diversity of real-world prediction targets. Unlike previous benchmarks, nearly one-third of the datasets in our BCCO benchmark contain missing values.
The BCCO-CLS benchmark comprises 106 datasets. The collection spans diverse sources, domains, and scales, and is designed to cover a wide range of problem characteristics, including the number of samples, number of features, number of classes, categorical–to-numerical feature ratio, sample-to-feature ratio, and proportion of missing values. This diversity allows for uniform binning along these dimensions, enabling results to be reported within bins and macro-averaged across bins, thereby providing a nearly unbiased assessment of model performance across heterogeneous task regimes. Additionally, datasets containing more than 50,000 training samples (The number of testing samples is not constrained), 10,000 features, or 10 target categories were excluded.
For each dataset in these benchmarks, we use the provided train-test split when available. If no predefined test set exists, the data are partitioned into a 70% training set and a 30% test set using stratified sampling to preserve the label distribution.
Baselines.
We compare LimiX with a range of state-of-the-art baseline models, categorized into tree-based models, neural networks (NN), and recent ICL-based approaches.
  • Tree-based approaches. We include XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), CatBoost (Dorogush et al., 2018), Random Forest (RF) (Breiman, 2001), and Extra Trees (ET) (Geurts et al., 2006). All models are optimized using the Optuna (Akiba et al., 2019) framework via 5-fold stratified cross-validation, with hyperparameters sampled from the ranges specified in Section A.2. Additionally, for AutoGluon-Tabular (Erickson et al., 2020), which automates workflows of model searching and ensemble, we use the default search space and set a default 600s time constraint for hyperparameter searching for each dataset.
  • NN-based approaches. We evaluate against SNN (Klambauer et al., 2017), AutoInt (Song et al., 2019), NODE (Popov et al., 2020), TabTransformer (Huang et al., 2020), GrowNet (Badirli et al., 2020), TabNet (Arik & Pfister, 2021), DCN-v2 (Wang et al., 2021), FT-Transformer (Gorishniy et al., 2021), MLP (Goodfellow et al., 2016; Gorishniy et al., 2021), ResNet (He et al., 2016; Gorishniy et al., 2021), SAINT (Somepalli et al., 2022), MLP-PLR (Gorishniy et al., 2022), DANets (Chen et al., 2022), TANGOS (Jeffares et al., 2023), T2G-Former (Yan et al., 2023), ExcelFormer (Chen et al., 2023b), Trompt (Chen et al., 2023c), TabR (Gorishniy et al., 2024), TabCaps (Chen et al., 2023a), RealMLP (Holzmüller et al., 2024), SwitchTab (Wu et al., 2024), and ModernNCA (Ye et al., 2025). These NN-based models are trained using the TALENT (Liu et al., 2024) Toolbox.
  • ICL-based models. Recent baselines includes TabPFN-v2 (Hollmann et al., 2025), TabICL (Qu et al., 2025), and Mitra (Zhang & Danielle, 2025).
Since TabDPT (Ma et al., 2024), another recent ICL-based model, is pretrained on real data that has a large overlap with datasets in the benchmarks, we do not adopt it as a baseline in the evaluation of standard classification and regression for a fair comparison. However, we include it in our experiments of fine-tuning in Section 7.6 to show the superiority of LimiX.
Metrics.
To evaluate model performance, we employ ROC AUC (area under the receiver operating characteristic curve), accuracy, and F1 score for classification tasks. For multi-class classification, the One-vs-One strategy is applied to both ROC AUC and F1 score calculations.
The critical difference diagram is also employed to compare the performance differences between LimiX and the baseline models. We conducted a Friedman test followed by a post-hoc Wilcoxon-Holm test, using a significance level of 0.05. In the diagram, the horizontal line indicates the range of ranks among which differences are not statistically significant.
Results.
Figure 7 shows that LimiX achieves the best performance among all methods on most datasets in BCCO-CLS. The most competitive baselines include other ICL-based models TabICL and TabPFN-v2, and the AutoML ensemble framework AutoGluon. In all critical difference diagrams, i.e. Figures 8, 9, 10, 11, 12 and 13, we can see that LimiX achieves the highest ranking in terms of all three evaluation metrics on BCCO-CLS, OpenML-CC18, and TALENT-CLS, where critical differences can be observed in most cases, indicating a significant margin.
For detailed quantitative results listed in Tables 1, 3, 5, 7, 9 and 11, LimiX outperforms all baselines on every benchmark in terms of both mean and rank of all three metrics. All these results clearly demonstrate that LimiX achieves state-of-the-art performance in terms of tabular classification tasks, surpassing not only traditional ensemble methods but also advanced ICL-based models.
Figure 7: The proportion of models achieving the best AUC. The length of each bar represents the proportion of the 106 datasets in BCCO-CLS where a given method achieves the highest AUC.
Subgroup analysis.
We use the sample subgroups when building BCCO-CLS to perform stratified analyses. Subgroups are defined based on the following criteria: the type of classification (binary or multi-class), the number of training samples, the ratio between the number of samples and features (length-to-width ratio), the proportion of categorical features, and the presence or absence of missing values. The number of training samples, categorical feature ratio, and length-to-width ratio are discretized into terciles (equal-frequency bins), ensuring that approximately the same number of datasets is allocated to each stratum and that the distribution of the dataset attribute which is not considered during analysis remains nearly uniform across strata.
From Figure 19, we observe that LimiX exhibits leading performance across all subgroups compared with other methods. Notably, after stratification, under some subgroups like the third subgroup in Figure 19(c) that indicates a larger training sample size, AutoGluon proves to be a strong competitor to ICL-based models. In these cases, LimiX is the only ICL-based model that outperforms AutoGluon. Notably, Figure 19(e) shows that, as the proportion of categorical features increases, performance for most baselines drops rapidly, reflecting the sparsity and high-cardinality challenges. In contrast, LimiX exhibits only modest degradation and remains comparatively stable across these regimes.
Table 1: Classification results on the BCCO-CLS benchmark. The best scores are shown in bold.
BCCO-CLS Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.871±0.008 0.804±0.010 0.731±0.020 2.547 2.151 2.764 TabICL 0.847±0.008 0.768±0.010 0.672±0.016 6.642 8.142 10.132 AutoGluon 0.846±0.000 0.771±0.001 0.677±0.001 7.811 7.934 9.302 TabPFN-v2 0.843±0.005 0.772±0.007 0.679±0.009 8.462 9.094 10.698 Mitra 0.836±0.017 0.764±0.018 0.664±0.035 10.104 11.038 12.849 XGBoost 0.834±0.005 0.762±0.007 0.674±0.012 9.887 10.208 10.887 LightGBM 0.832±0.005 0.763±0.006 0.678±0.011 10.802 10.075 10.698 RF 0.829±0.007 0.756±0.008 0.652±0.014 11.349 11.717 13.047 CatBoost 0.829±0.008 0.757±0.010 0.664±0.014 11.943 11.000 11.792 ET 0.825±0.005 0.745±0.006 0.618±0.014 12.406 13.925 16.783 ModernNCA 0.815±0.017 0.752±0.016 0.658±0.027 15.377 15.689 15.321 FT-Transformer 0.813±0.014 0.744±0.015 0.642±0.030 13.509 13.953 14.377 ExcelFormer 0.810±0.018 0.742±0.015 0.655±0.023 15.783 15.217 13.840 TabR 0.809±0.019 0.750±0.019 0.657±0.035 16.443 14.415 13.443 T2G-Former 0.808±0.021 0.742±0.020 0.646±0.034 14.500 14.491 13.670 MLP-PLR 0.804±0.007 0.733±0.011 0.635±0.018 15.717 16.179 15.377 ResNet 0.800±0.017 0.728±0.016 0.641±0.022 16.943 17.160 15.896 TANGOS 0.799±0.015 0.731±0.016 0.641±0.022 17.179 16.943 15.406 DCN-v2 0.794±0.020 0.725±0.018 0.618±0.035 16.953 16.925 16.698 RealMLP 0.794±0.020 0.738±0.014 0.644±0.021 18.255 16.462 15.208 SAINT 0.791±0.017 0.726±0.013 0.623±0.021 18.736 17.642 16.311 MLP 0.787±0.011 0.720±0.013 0.614±0.018 18.509 17.925 18.783 AutoInt 0.779±0.028 0.718±0.017 0.601±0.037 20.208 19.453 19.849 SNN 0.773±0.024 0.708±0.019 0.584±0.049 20.698 20.321 20.594 DANets 0.771±0.029 0.705±0.027 0.601±0.043 20.377 19.575 19.868 SwitchTab 0.766±0.021 0.700±0.018 0.590±0.036 21.396 21.991 20.868 TabTransformer 0.762±0.018 0.699±0.023 0.566±0.047 20.670 19.849 20.321 NODE 0.754±0.046 0.695±0.031 0.531±0.078 23.925 22.594 25.255 TabCaps 0.744±0.027 0.701±0.019 0.580±0.026 23.708 21.670 22.113 TabNet 0.712±0.047 0.685±0.028 0.561±0.041 27.151 24.660 24.774 GrowNet 0.682±0.033 0.641±0.028 0.522±0.040 27.726 27.368 26.387
Table 2: Statistical profile of the benchmark BCCO-CLS, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 4 12 48 23 33 1 259
# Classes 2 2 5 3 2 2 10
Missing Values (Ratio) 0 0 0.069 0.024 0.068 0 0.403
Categorical Features (Ratio) 0 0.384 0.929 0.399 0.337 0 1
Features w/ Missing Values (Ratio) 0 0 0.523 0.126 0.241 0 0.909
(a) AUC on the BCCO-CLS benchmark.
(b) Accuracy on the BCCO-CLS benchmark.
(c) F1-score on the BCCO-CLS benchmark.
Figure 8: Critical difference diagrams on BCCO-CLS benchmark.
Table 3: Classification results on the TALENT-CLS benchmark. The best scores are shown in bold.
TALENT-CLS Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.903±0.004 0.861±0.005 0.752±0.016 3.894 3.000 4.028 TabPFN-v2 0.895±0.002 0.850±0.003 0.727±0.008 6.056 5.860 7.670 TabICL 0.894±0.003 0.845±0.005 0.715±0.008 5.536 6.480 7.955 AutoGluon 0.891±0.000 0.845±0.000 0.719±0.001 6.229 6.642 7.425 Mitra 0.882±0.000 0.834±0.001 0.689±0.001 10.391 11.095 13.168 XGBoost 0.881±0.003 0.837±0.004 0.713±0.009 9.235 9.380 9.715 LightGBM 0.880±0.003 0.836±0.004 0.713±0.009 9.486 9.721 10.067 RF 0.877±0.003 0.828±0.004 0.691±0.010 12.307 12.793 13.927 CatBoost 0.876±0.005 0.828±0.006 0.704±0.014 11.453 11.497 11.743 ET 0.875±0.003 0.821±0.006 0.662±0.017 12.994 15.380 16.933 ExcelFormer 0.870±0.012 0.826±0.009 0.699±0.019 13.358 13.989 12.961 ResNet 0.866±0.009 0.825±0.006 0.695±0.014 14.933 14.721 13.207 MLP 0.862±0.008 0.817±0.008 0.675±0.017 16.419 16.140 16.346 TANGOS 0.861±0.009 0.818±0.007 0.684±0.020 16.011 15.978 14.754 ModernNCA 0.861±0.013 0.825±0.010 0.683±0.022 17.101 16.877 16.151 FT-Transformer 0.859±0.010 0.822±0.007 0.678±0.017 15.693 15.905 15.609 TabR 0.858±0.015 0.824±0.011 0.680±0.026 17.413 16.034 15.179 T2G-Former 0.858±0.016 0.823±0.010 0.683±0.019 15.777 15.536 15.134 DCN-v2 0.854±0.016 0.815±0.013 0.662±0.027 17.849 18.039 17.899 MLP-PLR 0.849±0.007 0.816±0.009 0.663±0.019 18.492 17.346 17.391 DANets 0.848±0.018 0.805±0.018 0.654±0.038 19.039 18.838 18.497 AutoInt 0.842±0.011 0.803±0.008 0.646±0.022 20.587 21.246 20.721 SwitchTab 0.842±0.013 0.795±0.011 0.637±0.029 21.346 21.894 21.089 RealMLP 0.839±0.017 0.820±0.007 0.678±0.017 20.212 17.687 15.994 SNN 0.836±0.020 0.796±0.016 0.625±0.041 22.011 21.955 22.089 TabCaps 0.834±0.014 0.813±0.008 0.654±0.017 21.324 18.223 18.687 TabTransformer 0.832±0.014 0.790±0.016 0.627±0.037 21.743 21.816 20.933 NODE 0.830±0.043 0.779±0.040 0.570±0.104 22.883 23.240 24.816 TabNet 0.818±0.035 0.794±0.021 0.630±0.041 25.006 22.603 23.039 SAINT 0.813±0.012 0.781±0.007 0.630±0.020 20.279 19.402 18.888 GrowNet 0.743±0.019 0.704±0.011 0.542±0.024 27.453 27.000 25.028
Table 4: Statistical profile of the benchmark TALENT-CLS, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 7 19 70 33 47 3 308
# Classes 2 2 10 6 14 2 100
Missing Values (Ratio) 0 0 0 0.001 0.008 0 0.1
Categorical Features (Ratio) 0 0.121 0.972 0.306 0.365 0 1
Features w/ Missing Values (Ratio) 0 0 0 0.025 0.121 0 0.979
(a) AUC on the TALENT-CLS benchmark.
(b) Accuracy on the TALENT-CLS benchmark.
(c) F1-score on the TALENT-CLS benchmark.
Figure 9: Critical difference diagram on the TALENT-CLS benchmark
Table 5: Classification results on the OpenML-CC18 benchmark. The best scores are shown in bold.
OpenML-cc18 Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.939±0.004 0.893±0.008 0.811±0.014 4.048 4.952 4.258 AutoGluon 0.932±0.000 0.885±0.001 0.790±0.003 5.823 6.694 7.048 TabPFN-v2 0.929±0.002 0.886±0.005 0.790±0.006 7.177 7.177 7.629 XGBoost 0.929±0.002 0.879±0.005 0.775±0.012 8.952 10.145 10.000 TabICL 0.927±0.004 0.875±0.008 0.782±0.011 8.242 9.677 9.613 LightGBM 0.927±0.003 0.879±0.005 0.775±0.015 9.323 9.532 9.565 CatBoost 0.926±0.005 0.870±0.008 0.770±0.016 11.258 12.726 12.613 RF 0.925±0.003 0.871±0.006 0.762±0.016 12.871 12.452 13.145 ET 0.922±0.003 0.861±0.006 0.721±0.017 12.903 15.016 16.839 Mitra 0.920±0.001 0.866±0.004 0.743±0.004 12.177 13.758 15.226 ExcelFormer 0.918±0.014 0.870±0.014 0.773±0.022 13.065 12.210 12.048 ResNet 0.913±0.009 0.860±0.011 0.764±0.019 13.435 13.613 12.806 TANGOS 0.910±0.007 0.863±0.010 0.759±0.025 13.355 13.339 13.484 MLP 0.908±0.007 0.857±0.008 0.743±0.019 13.887 13.984 15.016 T2G-Former 0.908±0.012 0.859±0.011 0.748±0.020 15.145 14.677 14.710 ModernNCA 0.906±0.009 0.858±0.013 0.747±0.021 16.661 16.129 15.532 FT-Transformer 0.904±0.015 0.856±0.012 0.739±0.019 16.532 15.306 16.194 TabR 0.900±0.015 0.863±0.011 0.757±0.025 16.661 13.484 13.113 DANets 0.900±0.021 0.840±0.029 0.715±0.053 18.145 17.613 18.581 DCN-v2 0.899±0.017 0.851±0.015 0.729±0.046 18.710 17.516 18.129 SNN 0.896±0.014 0.835±0.023 0.698±0.055 20.161 20.000 20.774 MLP-PLR 0.896±0.010 0.858±0.008 0.734±0.020 15.919 15.645 16.000 NODE 0.889±0.042 0.809±0.060 0.626±0.126 23.177 24.242 26.194 RealMLP 0.889±0.016 0.858±0.013 0.742±0.026 16.274 14.839 14.065 AutoInt 0.889±0.010 0.835±0.015 0.717±0.033 18.145 18.161 18.290 SwitchTab 0.881±0.031 0.819±0.039 0.671±0.065 21.403 20.790 21.516 TabCaps 0.875±0.024 0.848±0.017 0.692±0.042 21.597 18.613 20.161 TabNet 0.863±0.039 0.816±0.044 0.666±0.076 25.226 23.177 24.839 TabTransformer 0.829±0.021 0.776±0.022 0.616±0.051 22.565 21.645 22.387 GrowNet 0.816±0.027 0.748±0.027 0.581±0.050 25.403 25.855 24.935 SAINT 0.512±0.000 0.492±0.000 0.441±0.000 25.210 24.597 23.290
Table 6: Statistical profile of the benchmark OpenML-CC18, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 30 611 336 1344 4 10935
# Classes 2 3 10 6 7 2 46
Missing Values (Ratio) 0 0 0.002 0.004 0.019 0 0.139
Categorical Features (Ratio) 0 0.091 1 0.327 0.406 0 1
Features w/ Missing Values (Ratio) 0 0 0.204 0.054 0.165 0 0.757
(a) AUC on the OpenML-CC18 benchmark.
(b) Accuracy on the OpenML-CC18 benchmark.
(c) F1-score on the OpenML-CC18 benchmark.
Figure 10: Critical difference diagrams on the OpenML-CC18 benchmark.
Table 7: Classification results on the TabArena benchmark. The best scores are shown in bold.
TabArena Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.849±0.005 0.877±0.004 0.597±0.014 3.182 3.030 6.606 AutoGluon 0.844±0.001 0.870±0.000 0.574±0.001 5.212 4.636 8.364 LightGBM 0.841±0.002 0.868±0.002 0.574±0.011 6.758 7.667 9.848 TabICL 0.840±0.006 0.870±0.002 0.553±0.007 6.364 5.727 9.970 XGBoost 0.838±0.002 0.867±0.003 0.567±0.018 7.576 6.939 10.182 TabPFN-v2 0.838±0.002 0.872±0.002 0.589±0.004 6.545 4.000 8.242 RF 0.837±0.002 0.864±0.003 0.558±0.011 9.000 7.455 11.303 CatBoost 0.835±0.005 0.867±0.004 0.574±0.015 9.333 6.758 9.273 ET 0.833±0.003 0.857±0.003 0.505±0.019 10.030 10.333 15.515 Mitra 0.815±0.025 0.862±0.008 0.533±0.049 11.152 9.485 14.152 ExcelFormer 0.810±0.014 0.849±0.015 0.555±0.027 13.788 15.939 14.152 TANGOS 0.791±0.019 0.844±0.009 0.522±0.047 17.636 16.848 15.212 TabR 0.785±0.012 0.842±0.013 0.510±0.049 19.182 18.606 15.303 ModernNCA 0.783±0.020 0.846±0.006 0.511±0.025 18.636 19.424 16.152 ResNet 0.781±0.029 0.824±0.034 0.532±0.045 19.394 20.515 15.515 MLP-PLR 0.781±0.007 0.836±0.004 0.460±0.013 18.121 18.091 18.818 T2G-Former 0.779±0.037 0.822±0.047 0.482±0.092 17.727 19.091 15.061 RealMLP 0.777±0.020 0.822±0.015 0.512±0.032 19.788 20.030 13.697 MLP 0.772±0.016 0.822±0.014 0.459±0.019 17.576 16.697 18.727 FT-Transformer 0.770±0.054 0.803±0.067 0.468±0.089 19.545 20.667 17.788 DCN-v2 0.769±0.037 0.833±0.019 0.482±0.051 20.333 18.424 17.394 AutoInt 0.769±0.042 0.826±0.032 0.474±0.060 18.788 19.636 17.788 NODE 0.769±0.073 0.792±0.070 0.352±0.145 19.333 20.515 23.485 SNN 0.755±0.054 0.818±0.033 0.442±0.054 21.273 21.091 21.152 SwitchTab 0.754±0.040 0.799±0.030 0.409±0.082 22.303 23.939 21.485 DANets 0.749±0.050 0.776±0.084 0.453±0.073 22.182 22.848 17.545 TabCaps 0.742±0.019 0.837±0.007 0.471±0.020 21.394 17.000 18.818 TabTransformer 0.739±0.021 0.781±0.022 0.438±0.040 19.667 20.424 18.303 TabNet 0.709±0.089 0.789±0.070 0.438±0.081 24.848 21.364 22.303 SAINT 0.694±0.021 0.739±0.018 0.437±0.029 23.576 24.061 20.606 GrowNet 0.646±0.052 0.674±0.083 0.361±0.072 25.697 27.091 22.394
Table 8: Statistical profile of the benchmark TabArena, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 9 21 129 125 374 4 1776
# Classes 2 2 3 262 1577 2 9856
Missing Values (Ratio) 0 0 0.056 0.027 0.109 0 0.672
Categorical Features (Ratio) 0 0.47 1 0.45 0.366 0 1
Features w/ Missing Values (Ratio) 0 0 0.696 0.139 0.294 0 0.994
(a) AUC on the TabArena benchmark.
(b) Accuracy on the TabArena benchmark.
(c) F1-score on the TabArena benchmark.
Figure 11: Critical difference diagrams on the TabArena benchmark.
Table 9: Classification results on the PFN-CLS benchmark. The best scores are shown in bold.
PFN-CLS Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.923±0.005 0.862±0.009 0.786±0.015 2.034 3.000 2.862 TabPFN-v2 0.910±0.002 0.845±0.007 0.756±0.010 4.862 6.793 6.483 AutoGluon 0.906±0.001 0.835±0.001 0.738±0.002 5.172 7.069 6.793 TabICL 0.903±0.005 0.832±0.008 0.742±0.014 6.276 8.034 8.379 XGBoost 0.898±0.003 0.831±0.007 0.733±0.023 8.241 6.862 7.759 Mitra 0.897±0.010 0.826±0.017 0.719±0.032 8.621 9.552 11.276 RF 0.896±0.006 0.822±0.008 0.721±0.019 10.414 10.172 11.310 CatBoost 0.895±0.005 0.819±0.009 0.720±0.015 9.552 9.448 10.621 LightGBM 0.893±0.004 0.826±0.006 0.725±0.016 8.793 8.103 8.655 ET 0.893±0.003 0.809±0.007 0.675±0.018 11.345 12.759 14.586 ExcelFormer 0.883±0.008 0.812±0.012 0.713±0.023 13.379 11.621 12.897 ResNet 0.869±0.011 0.801±0.010 0.700±0.021 14.276 14.621 13.931 MLP 0.866±0.007 0.795±0.009 0.695±0.019 16.310 15.759 15.310 TANGOS 0.865±0.010 0.797±0.009 0.698±0.019 14.931 15.931 14.310 ModernNCA 0.860±0.011 0.798±0.010 0.692±0.020 17.345 15.759 14.724 SwitchTab 0.858±0.014 0.776±0.013 0.626±0.044 18.241 19.448 19.069 FT-Transformer 0.849±0.018 0.789±0.012 0.683±0.017 18.103 17.897 16.586 T2G-Former 0.848±0.024 0.792±0.016 0.688±0.031 17.379 15.897 15.897 MLP-PLR 0.848±0.006 0.786±0.007 0.650±0.020 18.414 18.828 19.931 DCN-v2 0.846±0.021 0.771±0.020 0.633±0.058 21.793 23.000 23.103 DANets 0.844±0.039 0.770±0.040 0.631±0.077 18.897 18.862 19.966 NODE 0.844±0.058 0.754±0.060 0.535±0.175 22.069 23.862 26.000 TabR 0.842±0.028 0.789±0.013 0.688±0.024 20.759 17.931 16.483 AutoInt 0.838±0.024 0.772±0.017 0.640±0.049 22.103 21.241 21.069 TabCaps 0.834±0.022 0.788±0.014 0.636±0.031 20.069 18.103 19.207 SNN 0.831±0.027 0.762±0.023 0.595±0.063 23.379 21.138 23.034 RealMLP 0.829±0.016 0.780±0.008 0.678±0.014 21.241 19.138 18.241 TabNet 0.825±0.043 0.768±0.027 0.631±0.059 24.759 22.586 23.793 TabTransformer 0.821±0.025 0.761±0.015 0.604±0.057 21.483 20.724 22.379 GrowNet 0.756±0.024 0.689±0.019 0.497±0.034 27.621 26.655 26.138 SAINT 0.708±0.033 0.669±0.020 0.563±0.037 22.241 21.379 20.276
Table 10: Statistical profile of the benchmark PFN-CLS, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 21 187 58 80 4 308
# Classes 2 2 7 4 2 2 10
Missing Values (Ratio) 0 0 0 0.001 0.006 0 0.032
Categorical Features (Ratio) 0 0.081 0.956 0.291 0.381 0 1
Features w/ Missing Values (Ratio) 0 0 0 0.016 0.086 0 0.474
(a) AUC on the PFN-CLS benchmark
(b) Accuracy on the PFN-CLS benchmark
(c) F1-score on the PFN-CLS benchmark
Figure 12: Critical difference diagrams on PFN-CLS benchmark
Table 11: Classification results on the TabZilla benchmark. The best scores are shown in bold.
TabZilla Mean Rank Model AUC ($\uparrow$) Acc. ($\uparrow$) F1 ($\uparrow$) AUC ($\downarrow$) Acc. ($\downarrow$) F1 ($\downarrow$) LimiX 0.943±0.005 0.885±0.018 0.836±0.022 4.556 4.593 5.741 AutoGluon 0.933±0.001 0.871±0.002 0.803±0.008 6.778 7.148 8.593 TabICL 0.933±0.006 0.864±0.016 0.803±0.023 8.444 9.185 10.148 TabPFN-v2 0.929±0.002 0.863±0.008 0.797±0.012 7.333 8.481 8.889 XGBoost 0.929±0.003 0.863±0.005 0.789±0.011 8.741 10.074 11.185 LightGBM 0.927±0.003 0.863±0.007 0.796±0.010 10.444 9.667 10.185 RF 0.924±0.004 0.852±0.006 0.773±0.012 10.926 11.852 12.481 CatBoost 0.922±0.006 0.848±0.011 0.780±0.020 12.185 13.296 13.185 Mitra 0.915±0.010 0.841±0.015 0.758±0.037 14.074 14.074 15.037 ExcelFormer 0.915±0.012 0.861±0.014 0.802±0.029 13.815 11.444 11.444 ET 0.912±0.004 0.837±0.008 0.745±0.014 14.407 14.926 16.222 T2G-Former 0.909±0.009 0.852±0.012 0.790±0.014 14.148 14.333 14.407 TANGOS 0.909±0.006 0.841±0.014 0.776±0.032 15.148 13.704 14.148 ResNet 0.908±0.008 0.834±0.017 0.769±0.028 16.444 15.148 14.741 ModernNCA 0.907±0.007 0.850±0.012 0.794±0.017 14.222 14.556 13.926 MLP-PLR 0.906±0.005 0.847±0.008 0.773±0.016 14.963 14.889 15.926 TabR 0.904±0.014 0.853±0.014 0.793±0.029 14.963 13.185 13.630 DCN-v2 0.904±0.008 0.844±0.013 0.781±0.018 17.556 16.519 17.111 MLP 0.903±0.013 0.825±0.021 0.747±0.042 16.148 16.556 17.000 FT-Transformer 0.903±0.023 0.842±0.027 0.769±0.058 15.926 15.926 15.370 AutoInt 0.896±0.020 0.833±0.030 0.748±0.069 16.704 16.259 16.519 RealMLP 0.892±0.012 0.846±0.009 0.781±0.018 19.370 16.407 15.741 TabCaps 0.887±0.014 0.816±0.016 0.729±0.032 22.926 20.296 20.630 DANets 0.881±0.039 0.800±0.050 0.712±0.091 20.889 19.963 20.556 SNN 0.874±0.021 0.816±0.020 0.706±0.038 20.444 18.704 19.741 NODE 0.869±0.046 0.784±0.046 0.633±0.101 23.481 23.593 25.556 SwitchTab 0.860±0.030 0.764±0.045 0.660±0.073 21.889 23.519 23.148 TabNet 0.860±0.035 0.771±0.051 0.668±0.070 26.407 25.778 25.407 GrowNet 0.829±0.019 0.732±0.017 0.651±0.032 25.444 25.037 24.037 SAINT 0.824±0.013 0.764±0.017 0.680±0.046 19.148 18.852 18.556 TabTransformer 0.814±0.020 0.759±0.024 0.659±0.051 21.333 19.852 20.778
Table 12: Statistical profile of the benchmark TabZilla, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 24 132 226 756 4 4296
# Classes 2 2 10 7 16 2 100
Missing Values (Ratio) 0 0 0.078 0.022 0.059 0 0.205
Categorical Features (Ratio) 0 0.528 1 0.474 0.39 0 1
Features w/ Missing Values (Ratio) 0 0 0.483 0.102 0.218 0 0.808
(a) AUC on the TabZilla benchmark.
(b) Accuracy on the TabZilla benchmark.
(c) F1-score on the TabZilla benchmark.
Figure 13: Critical difference diagrams on the TabZilla benchmark.
7.2 Regression
Benchmark.
For the quantitative evaluation of regression performance, we leverage four benchmarks, including three open-source benchmark, TALENT-REG (Liu et al., 2024), PFN-REG (Hollmann et al., 2025), and CTR23 (Fischer et al., 2023). After applying filters to exclude datasets containing more than 50,000 training samples or 10,000 features, we have 33 datasets from CTR23, 28 from PFN-REG, and 99 from TALENT-REG.
Similar to BCCO-CLS, we introduce BCCO-REG, a balanced regression benchmark comprising 50 datasets. For each dataset, we adopt the provided train-test split when available. If no predefined test set is available, we partition the data into a 70% training set and a 30% testing set randomly. Similar to the protocol employed in classification tasks, regression datasets with more than 50,000 training samples or 10,000 features are excluded.
Baselines.
For comparison, we include tree-based methods, NN-based methods, AutoML frameworks, and ICL-based models, which are basically consistent with baselines employed in classification tasks described above. TabICL and TabCaps are excluded because it cannot handle regression tasks. We add another baseline, DNNR (Nader et al., 2022), which is specifically designed for regression. The model training procedure remains aligned with that employed in the classification experiments.
Metrics.
To evaluate regression performance properly, we employ normalized RMSE and R2 as the two evaluation metrics. We also analyze the ranks of models with respect to both metrics.
Figure 14: The proportion of models achieving the best R2. The length of each bar represents the proportion of 106 datasets in which a given method achieved the highest R2 on the BCCO-REG benchmark.
Results.
As shown by critical difference diagrams, i.e. Figures 15, 16, 17 and 18, LimiX always achieves the best performance, outperforming strong baselines like AutoGluon, XGBoost, and TabPFN-v2. Quantitative results in Tables 13, 19, 15 and 17 also confirms that LimiX achieves state-of-the-art regression performance, leading in both normalized RMSE and R2.
Subgroup analysis.
Similar to the subgroup analysis in classification, for regression, we also use the sample subgroups when building BCCO-REG to perform stratified analyses, whose criteria include the number of training samples, the ratio between the number of samples and features (length-to-width ratio), and the proportion of categorical features. Figure 20 shows that LimiX outperforms other methods across all subgroups compared with other methods. For Figures 20(b), 20(c) and 20(d), we can see that performance of Mitra substantially drops for the second subgroup, while LimiX do not exhibit a large drop.
(a) R2 on the BCCO-REG benchmark
(b) RMSE on the BCCO-REG benchmark
Figure 15: Critical difference diagram on the BCCO-REG benchmark
(a) R2 on the TALENT-REG benchmark
(b) RMSE on TALENT-REG benchmark
Figure 16: Critical difference diagram on the TALENT-REG benchmark
(a) R2 on the CTR23 benchmark
(b) RMSE on the CTR23 benchmark
Figure 17: Critical difference diagram on the CTR23 benchmark
(a) R2 on the PFN-REG benchmark
(b) RMSE on PFN-REG benchmark
Figure 18: Critical difference diagram on the PFN-REG benchmark
Table 13: Regression results on the BCCO-REG benchmark. The best scores are shown in bold.
BCCO-REG
Mean Rank
Model R2 ($\uparrow$) RMSE ($\downarrow$) R2 ($\downarrow$) RMSE ($\downarrow$)
LimiX 0.794±0.004 0.386±0.004 3.100 3.760
AutoGluon 0.781±0.001 0.398±0.000 3.700 3.660
TabPFN-v2 0.772±0.001 0.404±0.002 5.000 4.900
XGBoost 0.764±0.006 0.415±0.005 7.740 7.780
ET 0.757±0.005 0.431±0.006 9.960 9.900
RF 0.752±0.007 0.438±0.009 11.260 11.220
T2G-Former 0.743±0.011 0.442±0.012 12.860 12.820
ExcelFormer 0.743±0.015 0.443±0.012 11.460 11.520
CatBoost 0.741±0.026 0.427±0.020 10.520 10.360
DCN-v2 0.739±0.015 0.448±0.014 13.780 13.780
FT-Transformer 0.737±0.009 0.448±0.010 13.440 13.420
MLP-PLR 0.734±0.012 0.453±0.009 15.720 15.720
TabR 0.733±0.017 0.448±0.011 14.460 14.520
RealMLP 0.725±0.029 0.441±0.010 12.420 12.380
AutoInt 0.724±0.015 0.465±0.018 17.220 17.140
ResNet 0.720±0.009 0.468±0.010 15.020 14.980
TANGOS 0.719±0.010 0.468±0.009 15.680 15.680
LightGBM 0.715±0.019 0.423±0.005 7.980 7.940
MLP 0.701±0.013 0.487±0.010 16.940 17.000
SAINT 0.701±0.022 0.481±0.019 16.520 16.540
TabNet 0.670±0.024 0.516±0.024 20.640 20.580
Mitra 0.667±0.027 0.474±0.015 15.700 15.620
NODE 0.643±0.099 0.543±0.097 20.100 20.020
ModernNCA 0.598±0.195 0.471±0.038 15.180 15.080
SNN 0.434±0.111 0.720±0.072 25.200 25.200
GrowNet 0.201±0.182 0.864±0.110 26.840 26.820
DANets 0.005±0.008 0.979±0.004 28.440 28.480
SwitchTab 0.001±0.032 0.981±0.016 28.700 28.700
TabTransformer 0.000±0.016 0.981±0.008 28.600 28.640
DNNR -2.152±0.150 1.329±0.027 20.820 20.840
Table 14: Statistical profile of the benchmark BCCO-REG, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 12 33 16 14 4 81
Missing Values (Ratio) 0 0 0 0 0 0 0.001
Categorical Features (Ratio) 0 0.16 0.65 0.242 0.264 0 0.967
Features w/ Missing Values (Ratio) 0 0 0 0.008 0.051 0 0.333
Table 15: Regression results on the TALENT-REG benchmark. The best scores are shown in bold.
TALENT-REG
Mean Rank
Model R2 ($\uparrow$) RMSE ($\downarrow$) R2 ($\downarrow$) RMSE ($\downarrow$)
LimiX 0.735±0.009 0.433±0.009 2.667 3.162
AutoGluon 0.722±0.006 0.448±0.009 4.313 4.414
XGBoost 0.710±0.006 0.462±0.005 6.768 6.697
LightGBM 0.707±0.004 0.461±0.004 7.485 7.455
CatBoost 0.700±0.015 0.471±0.018 9.667 9.606
RF 0.697±0.008 0.474±0.009 9.788 9.788
ET 0.696±0.008 0.476±0.007 9.242 9.192
TabPFN-v2 0.695±0.003 0.465±0.002 5.667 5.576
RealMLP 0.656±0.013 0.510±0.010 12.636 12.636
T2G-Former 0.656±0.019 0.512±0.017 13.364 13.303
MLP-PLR 0.653±0.017 0.521±0.017 15.242 15.202
ExcelFormer 0.653±0.016 0.512±0.012 12.444 12.485
TabR 0.651±0.023 0.516±0.021 14.434 14.394
FT-Transformer 0.648±0.019 0.519±0.018 15.000 14.949
AutoInt 0.636±0.020 0.538±0.021 18.152 18.162
ModernNCA 0.633±0.033 0.530±0.029 15.758 15.747
Mitra 0.602±0.053 0.547±0.022 14.909 14.899
TANGOS 0.592±0.027 0.547±0.013 16.202 16.172
TabNet 0.576±0.042 0.586±0.032 20.798 20.737
NODE 0.568±0.096 0.600±0.091 19.636 19.495
ResNet 0.562±0.095 0.550±0.023 15.697 15.667
MLP 0.556±0.025 0.564±0.008 16.677 16.717
SNN 0.344±0.138 0.777±0.069 24.485 24.505
DANets 0.005±0.009 0.998±0.004 27.475 27.495
TabTransformer 0.001±0.017 1.001±0.008 27.606 27.657
SwitchTab -0.002±0.027 1.002±0.013 27.980 28.010
GrowNet -0.182±0.319 0.920±0.120 26.162 26.192
DCN-v2 -0.361±0.161 0.818±0.182 14.717 14.747
SAINT -1.541±0.014 0.571±0.010 17.495 17.364
DNNR -9.172±5.278 2.528±0.814 22.525 22.566
Table 16: Statistical profile of the benchmark TALENT-REG, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 11 79 30 49 4 266
Missing Values (Ratio) 0 0 0 0.007 0.034 0 0.283
Categorical Features (Ratio) 0 0.131 0.667 0.238 0.283 0 1
Features w/ Missing Values (Ratio) 0 0 0.013 0.043 0.161 0 0.875
Table 17: Regression results on the CTR23 benchmark. The best scores are shown in bold.
CTR23
Mean Rank
Model R2 ($\uparrow$) RMSE ($\downarrow$) R2 ($\downarrow$) RMSE ($\downarrow$)
LimiX 0.745±0.008 0.477±0.007 3.879 3.970
AutoGluon 0.725±0.001 0.497±0.001 4.818 4.727
TabPFN-v2 0.716±0.002 0.503±0.002 7.182 7.061
XGBoost 0.712±0.008 0.511±0.007 8.000 8.061
LightGBM 0.706±0.004 0.516±0.005 9.000 9.091
CatBoost 0.700±0.022 0.528±0.030 10.455 10.455
ET 0.697±0.007 0.535±0.008 9.758 9.818
RF 0.694±0.010 0.539±0.012 10.667 10.667
T2G-Former 0.674±0.013 0.544±0.014 13.061 13.030
MLP-PLR 0.672±0.013 0.553±0.014 15.000 15.030
TabR 0.671±0.014 0.543±0.014 14.182 14.121
DCN-v2 0.670±0.015 0.545±0.015 13.788 13.848
FT-Transformer 0.667±0.011 0.549±0.011 14.061 14.152
ModernNCA 0.667±0.027 0.550±0.027 15.121 14.788
ExcelFormer 0.665±0.014 0.556±0.013 11.727 11.818
RealMLP 0.661±0.013 0.549±0.011 13.788 13.818
AutoInt 0.655±0.015 0.568±0.018 16.970 16.970
SAINT 0.654±0.015 0.561±0.017 15.970 15.879
ResNet 0.645±0.018 0.587±0.017 13.667 13.727
TANGOS 0.642±0.014 0.586±0.013 15.818 15.848
Mitra 0.624±0.017 0.583±0.016 15.788 15.909
MLP 0.608±0.008 0.623±0.008 15.000 15.000
TabNet 0.605±0.046 0.623±0.041 19.606 19.667
NODE 0.568±0.105 0.666±0.096 19.636 19.485
SNN 0.369±0.073 0.834±0.048 24.727 24.727
GrowNet 0.185±0.184 0.944±0.113 26.606 26.576
DANets 0.001±0.010 1.052±0.005 28.000 28.000
TabTransformer 0.000±0.016 1.053±0.008 28.000 28.000
SwitchTab -0.006±0.025 1.057±0.013 29.061 29.061
DNNR -2.969±0.262 1.651±0.040 21.667 21.697
Table 18: Statistical profile of the benchmark CTR23, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 11 39 20 22 5 116
Missing Values (Ratio) 0 0 0 0.008 0.036 0 0.209
Categorical Features (Ratio) 0 0.143 0.842 0.312 0.345 0 1
Features w/ Missing Values (Ratio) 0 0 0 0.011 0.05 0 0.286
Table 19: Regression results on the PFN-REG benchmark. The best scores are shown in bold.
PFN-REG
Mean Rank
Model R2 ($\uparrow$) RMSE ($\downarrow$) R2 ($\downarrow$) RMSE ($\downarrow$)
LimiX 0.692±0.012 0.461±0.011 4.143 4.964
AutoGluon 0.677±0.000 0.482±0.001 6.607 6.607
TabPFN-v2 0.676±0.004 0.475±0.002 7.250 7.036
XGBoost 0.671±0.007 0.487±0.007 8.679 8.679
RealMLP 0.664±0.012 0.484±0.010 9.429 9.357
LightGBM 0.664±0.008 0.496±0.008 9.714 9.643
CatBoost 0.661±0.014 0.498±0.014 10.607 10.571
ET 0.652±0.007 0.515±0.008 10.643 10.571
ExcelFormer 0.646±0.013 0.513±0.013 13.179 13.214
ModernNCA 0.645±0.021 0.508±0.015 14.286 14.464
RF 0.643±0.010 0.524±0.011 12.929 13.000
T2G-Former 0.640±0.027 0.503±0.022 13.143 13.036
MLP-PLR 0.638±0.017 0.513±0.019 14.429 14.500
DCN-v2 0.638±0.024 0.511±0.022 14.786 14.607
TabR 0.636±0.036 0.509±0.030 14.500 14.357
FT-Transformer 0.636±0.033 0.509±0.027 14.107 13.964
Mitra 0.630±0.014 0.531±0.016 12.821 12.857
AutoInt 0.610±0.031 0.539±0.033 18.214 18.107
ResNet 0.599±0.017 0.553±0.013 14.393 14.429
TANGOS 0.587±0.024 0.562±0.025 14.679 14.821
MLP 0.577±0.023 0.576±0.016 14.107 14.250
NODE 0.496±0.135 0.666±0.117 20.071 19.893
TabNet 0.430±0.216 0.643±0.106 21.107 21.036
SNN 0.378±0.079 0.756±0.056 24.179 24.214
GrowNet 0.096±0.141 0.942±0.080 25.679 25.643
DANets 0.001±0.012 0.997±0.006 26.893 26.929
TabTransformer -0.021±0.029 1.007±0.014 28.357 28.393
SwitchTab -0.026±0.034 1.011±0.017 27.429 27.464
DNNR -7.850±0.583 2.224±1.125 22.286 22.429
SAINT -8.088±0.020 0.612±0.014 16.357 15.964
Table 20: Statistical profile of the benchmark PFN-REG, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 16 176 66 93 3 376
Missing Values (Ratio) 0 0 0.089 0.029 0.076 0 0.335
Categorical Features (Ratio) 0 0.209 0.949 0.355 0.381 0 1
Features w/ Missing Values (Ratio) 0 0 0.211 0.059 0.169 0 0.841
(a) All datasets
(b) Binary classification
(c) Length-to-width ratio
(d) Presence of missing values
(e) Categorical-feature proportion
(f) Number of training samples
Figure 19: AUC on subsets with various sample size, number of classes, categorical–to-numerical feature ratios, missing values, and sample-to-feature ratios.
(a) All datasets
(b) Length-to-width ratio
(c) Categorical-feature proportion
(d) Number of training samples
Figure 20: R2 on subsets with various sample size, categorical–to-numerical feature ratios, and sample-to-feature ratios.
7.3 Missing Value Imputation
The ubiquity of missing values in tabular data is harmful to downstream tasks or statistical analyses (Lin & Tsai, 2020). Meanwhile, the problem of missing value imputation poses a bigger challenge compared with standard classification and regression tasks since it requires modeling of the joint distribution of $P(X,Y)$ instead of $P(Y|X)$ only. A noteworthy capability of LimiX is missing value imputation, which naturally stems from the mask modeling during pretraining. Although there are existing deep learning approaches for missing value imputation (Jarrett et al., 2022; Zhang et al., 2025), they all require additional training on unseen datasets. In contrast, LimiX is the first model to be capable of missing value imputation on unseen datasets via in-context learning without additional training. This brings great convenience to the usage in downstream tasks.
We conduct experiments on the following datasets: Analcatdata BroadwayMult (Simonoff, 2003; Vanschoren, 2014a), Early Stage Diabetes 2020 (Islam et al., 2019; 2020), Forty Soybean Cultivars from Subsequent Harvests (de Oliveira et al., 2023; Rodrigues de Oliveira & Mario Zuffo, 2023), HCC survival (Santos et al., 2015b; a), Vehicle (Siebert, 1987; Mowforth & Shepherd, 1987), Eucalyptus (van Rijn, 2014), and Z-Alizadeh Sani (Alizadehsani et al., 2013). To simulate the scenario of missing value imputation, we randomly mask a fixed proportion $\alpha$ of cells in ${\mathbf{x}}_{te}$ and try to recover values of these cells. In our experiments, we set $\alpha=0.05$. Baselines include: Mean/Mode (use mean for continuous features and mode for categorical features), k-nearest neighbors (KNN), MICE (Van Buuren & Groothuis-Oudshoorn, 2011), MissForest (Stekhoven & Bühlmann, 2012), GAIN (Yoon et al., 2018), MIWAE (Mattei & Frellsen, 2019), and HyperImpute (Jarrett et al., 2022). For continuous features, we normalize each feature using MinMaxScaler and calculate RMSE between imputed values and ground truth values after normalization as the evaluation metric. For categorical features, the error rate is used as the evaluation metric. Table 21 shows that LimiX consistently outperforms previous baselines, all of which require additional training or fitting.
Table 21: Evaluation of missing value imputation. For continuous features, we calculate RMSE. For categorical features, we calculate classification error. Our method outperforms previous missing value imputation methods.
Metric Regression RMSE (↓) Classification Error (↓)
Method Analcatdata Early Diabetes Harvests Hcc Survival Vehicle Eucalyptus Z-Alizadeh Sani
Mean/Mode 0.321 0.235 0.179 0.351 0.209 0.804 0.200
KNN 0.358 0.205 0.158 0.246 0.083 0.275 0.206
MICE 0.294 0.244 0.155 0.234 0.102 0.627 0.181
MissForest 0.203 0.223 0.136 0.215 0.107 0.137 0.156
GAIN 0.299 0.328 0.196 0.413 0.102 0.647 0.175
MIWAE 0.561 0.478 0.322 0.639 0.415 0.706 0.294
HyperImpute 0.272 0.270 0.152 0.297 0.086 0.647 0.194
LimiX 0.194 0.161 0.104 0.188 0.064 0.118 0.131
7.4 Robustness
Neural networks have been found to be vulnerable to various types of perturbations and attacks (Carlini & Wagner, 2017; Akhtar & Mian, 2018). In order to investigate the robustness of LimiX, following Hollmann et al. (2025), we conduct experiments under two types of controlled perturbations: adding uninformative features and outliers.
For uninformative features, we randomly select columns from the original dataset, shuffle each column, and concatenate them to the original dataset. For outliers, we multiply an outlier coefficient to each cell value in the original dataset with a probability of 2%. The outlier coefficient is randomly chosen between 0 and the outlier factor.
In Figure 21, we show Normalized AUC in classification tasks under the two types of perturbations. We compare LimiX with TabPFN-v2, TabICL, and CatBoost. The left figure shows that even when adding up to $90\%$ uninformative features, the normalized AUC of LimiX remains nearly unchanged. In contrast, the performance of TabICL and CatBoost drops significantly. This indicates that LimiX is much more robust than TabICL and CatBoost. The right figure shows that LimiX consistently outperforms other methods regardless of the outlier factor. In Figure 22, we show RMSE in regression tasks under perturbations. In the left figure of adding uninformative features, a similar phenomenon is observed as that in classification tasks. In the right figure of adding outliers, we find that RMSE of TabPFN-v2 rapidly increases when the outlier factor grows from 100 to 10000, while LimiX do not. This demonstrates the superior robustness of LimiX compared with TabPFN-v2 in regression tasks. Overall, LimiX exhibits superior robustness relative to the baselines.
Figure 21: Robustness analysis in classification Tasks. LimiX consistently exhibits the best performance and superior robustness under perturbations.
Figure 22: Robustness analysis in regression tasks. LimiX consistently exhibits the best performance and superior robustness under perturbations.
7.5 Embedding
Pretrained models (Devlin et al., 2019; Oquab et al., 2024) are generally expected to provide effective and transferable feature representations that can be leveraged for downstream tasks. To assess the quality of embeddings extracted by LimiX, we conduct experiments on six datasets of various sample sizes and numbers of categories. We compare embeddings of LimiX with those of MLP (Goodfellow et al., 2016; Gorishniy et al., 2021), ResNet (He et al., 2016; Gorishniy et al., 2021), ModernNCA (Ye et al., 2025), TabPFN-v2 (Hollmann et al., 2025), and TabICL (Qu et al., 2025). For each model, we treat representations prior to the classification head as embeddings. Figure 23 shows t-SNE visualization of embeddings. We can see that embeddings of different categories extracted by LimiX are more separated than those extracted by other models.
To further evaluate the quality of embeddings extracted by LimiX, we additionally conduct linear probing experiments, which are widely adopted in the analysis of feature representations (Kumar et al., 2022). The experiments are conducted on BCCO-CLS. Table 22 shows that embeddings of LimiX achieve the highest average AUC and rank among the three ICL-based models. Overall, LimiX consistently outperforms baselines in both qualitative and quantitative experiments.
Figure 23: t-SNE visualization of embeddings. Different colors represent different categories. Embeddings extracted by LimiX are better separated between categories than other models.
Table 22: Results of linear probing on BCCO-CLS. LimiX outperforms the other two ICL-based models, indicating the strongest embedding capability.
Model AUC(↑) Rank(↓)
TabPFN-v2 0.832 2.189
TabICL 0.838 1.981
LimiX 0.850 1.792
7.6 Fine-tuning
Long contexts increase memory cost and can degrade optimization during fine-tuning. Therefore, it is common practice to subsample the training corpus to a manageable budget, usually via random or KNN selection (Thomas et al., 2024; Xu et al., 2025). Based on our attention-based retrieval strategy, we adopt a retrieval-guided downsampling scheme that concentrates on locally most relevant patterns within each dataset, thus improving sample efficiency and prediction performance.
For each dataset, we split the original training set into a retrieval pool and a query set. For each instance in the query set, we select samples from the retrieval pool via the strategy proposed in Section 5 as in-context samples and construct an episode of in-context learning. Fine-tuning then proceeds over these episodes rather than over full, unfiltered contexts. This strategy substantially reduces the number of epochs required to achieve good performance. Note that the retrieved in-context sets may contain duplicated samples, which leads to a trade-off between relevance and diversity. We empirically find that the retrieved context size controls the balance between computation efficiency and prediction performance.
We compare LimiX with some representative baselines that are trained or fine-tuned on real datasets, including TabDPT (Ma et al., 2024). For the fine-tuning of TabPFN-v2, we adopt the strategy and hyperparameter configurations from its official repository [^2] https://github.com/PriorLabs/TabPFN. As shown in Tables 23 and 24, LimiX significantly outperforms the others across various metrics in most benchmarks after fine-tuning. Meanwhile, we illustrate the change of AUC for LimiX on some datasets of BCCO-CLS before and after fine-tuning. From Figure 24, a significant improvement can be observed in most cases.
Figure 24: The improvement of fine-tuning in AUC on the BCCO-CLS benchmark.
Table 23: Performance comparison of representative methods on benchmarks of classification. "FT" indicates the model performance after fine-tuning. TabDPT-NE represents the performance of TabDPT without ensemble while its default version has an ensemble size of 8. The best scores are shown in bold.
Benchmark Method AUC (↑) Acc. (↑) F1 (↑)
AUC-Rank
Acc.-Rank
F1-Rank
BCCO-CLS XGBoost 0.834 0.762 0.674 5.811 5.660 5.509
CatBoost 0.829 0.757 0.664 6.264 6.000 5.802
TabDPT-NE 0.841 0.774 0.685 5.085 4.679 4.783
TabDPT 0.846 0.777 0.687 4.453 4.358 4.604
TabPFN-v2 0.843 0.772 0.679 5.170 5.208 5.274
TabPFN-v2-FT 0.842 0.773 0.678 5.198 4.925 5.330
LimiX 0.871 0.804 0.731 2.538 2.575 2.651
LimiX-FT 0.873 0.806 0.733 1.396 1.434 1.481
TALENT-CLS XGBoost 0.881 0.837 0.713 5.603 5.564 5.402
CatBoost 0.876 0.828 0.704 6.318 6.229 6.000
TabDPT-NE 0.891 0.846 0.719 4.955 4.754 4.933
TabDPT 0.893 0.849 0.723 4.179 4.207 4.413
TabPFN-v2 0.895 0.850 0.727 4.436 4.464 4.570
TabPFN-v2-FT 0.889 0.842 0.718 5.212 5.156 5.106
LimiX 0.903 0.861 0.752 3.061 2.860 2.933
LimiX-FT 0.904 0.863 0.755 1.726 1.598 1.777
OpenML-CC18 XGBoost 0.929 0.879 0.775 5.177 5.194 5.500
CatBoost 0.926 0.870 0.770 6.548 6.032 6.274
TabDPT-NE 0.927 0.881 0.786 4.677 4.758 4.984
TabDPT 0.930 0.885 0.799 3.919 3.919 4.161
TabPFN-v2 0.929 0.886 0.790 5.065 4.468 4.613
TabPFN-v2-FT 0.933 0.889 0.883 4.161 3.823 3.032
LimiX 0.939 0.893 0.811 2.952 3.677 3.613
LimiX-FT 0.941 0.894 0.813 1.500 2.306 2.435
TabArena XGBoost 0.838 0.867 0.567 5.424 6.030 5.788
CatBoost 0.835 0.867 0.574 6.242 5.545 5.121
TabDPT-NE 0.891 0.846 0.719 4.950 4.749 4.927
TabDPT 0.839 0.870 0.580 4.970 4.636 4.667
TabPFN-v2 0.838 0.872 0.589 4.939 4.303 4.818
TabPFN-v2-FT 0.850 0.874 0.713 3.697 4.273 2.682
LimiX 0.849 0.877 0.597 3.242 3.030 4.212
LimiX-FT 0.851 0.878 0.600 2.030 1.697 2.939
TabZilla XGBoost 0.929 0.863 0.789 5.185 5.148 5.481
CatBoost 0.922 0.848 0.780 6.556 6.037 6.074
TabDPT-NE 0.932 0.880 0.824 4.333 3.963 4.037
TabDPT 0.934 0.882 0.824 3.815 3.815 4.148
TabPFN-v2 0.929 0.863 0.797 5.111 4.815 5.185
TabPFN-v2-FT 0.932 0.876 0.847 4.185 4.556 3.318
LimiX 0.943 0.885 0.836 3.556 3.444 3.889
LimiX-FT 0.944 0.887 0.838 2.074 2.407 3.136
PFN-CLS XGBoost 0.898 0.831 0.733 6.000 5.138 5.103
CatBoost 0.895 0.819 0.720 6.310 5.828 6.138
TabDPT-NE 0.891 0.822 0.727 5.034 5.483 5.828
TabDPT 0.896 0.833 0.742 4.655 4.103 4.448
TabPFN-v2 0.910 0.845 0.756 4.138 4.655 4.552
TabPFN-v2-FT 0.900 0.837 0.759 5.379 5.310 4.828
LimiX 0.923 0.862 0.786 2.241 2.690 2.759
LimiX-FT 0.924 0.864 0.788 1.276 1.517 1.724
Table 24: Performance comparison of representative methods on benchmarks of regression. "FT" indicates the model performance after fine-tuning. TabDPT-NE represents the performance of TabDPT without ensemble while its default version has an ensemble size of 8. The best scores are shown in bold.
Benchmark Method R2 (↑) RMSE (↓) R2-Rank (↓) RMSE-Rank (↓)
BCCO-REG XGBoost 0.764 0.415 5.820 5.740
CatBoost 0.741 0.427 6.600 6.460
TabDPT-NE 0.769 0.410 5.220 5.060
TabDPT 0.772 0.406 4.460 4.260
TabPFN-v2 0.772 0.404 4.600 4.280
TabPFN-v2-FT 0.777 0.399 4.020 3.720
LimiX 0.794 0.386 3.320 2.860
LimiX-FT 0.796 0.385 1.960 2.880
TALENT-REG XGBoost 0.710 0.462 5.545 5.404
CatBoost 0.700 0.471 6.485 6.434
TabDPT-NE 0.709 0.461 5.061 4.929
TabDPT 0.711 0.458 4.263 4.081
TabPFN-v2 0.695 0.465 4.980 4.747
TabPFN-v2-FT 0.702 0.459 4.879 4.657
LimiX 0.735 0.433 2.970 2.566
LimiX-FT 0.737 0.417 1.818 2.177
CTR23 XGBoost 0.712 0.511 5.636 5.606
CatBoost 0.700 0.528 6.273 6.212
TabDPT-NE 0.728 0.500 4.909 4.818
TabDPT 0.731 0.498 4.273 4.121
TabPFN-v2 0.716 0.503 4.545 4.394
TabPFN-v2-FT 0.722 0.498 4.636 4.545
LimiX 0.745 0.477 3.576 2.818
LimiX-FT 0.748 0.473 2.152 2.437
PFN-REG XGBoost 0.671 0.487 4.929 4.714
CatBoost 0.661 0.498 6.214 6.071
TabDPT-NE 0.672 0.491 5.393 5.071
TabDPT 0.675 0.484 4.607 4.500
TabPFN-v2 0.676 0.475 4.393 4.143
TabPFN-v2-FT 0.687 0.466 4.179 3.857
LimiX 0.692 0.461 3.857 3.571
LimiX-FT 0.695 0.466 2.429 4.071
7.7 Data Generation
Data generation is a challenging and meaningful task across multiple modalities since it needs to capture the joint distribution of $P(X,Y)$ compared with prediction tasks that focus on modeling $P(Y|X)$. While there is abundant literature on the generation of images (Croitoru et al., 2023), videos (Xing et al., 2024), and text (Li et al., 2024b), relatively less attention has been paid to tabular data generation. However, data generation is even more critical for tabular data due to its scarcity and possible privacy issues (Hernandez et al., 2022). As a foundation model, LimiX has the ability of tabular data generation given an unseen real dataset. A significant advantage is that it does not require additional training of generative models. Following TabPFN-v2 (Hollmann et al., 2025), firstly we conduct data generation in an iterative style. For the first column, we sample from the empirical categorical distribution if it is a categorical feature, or we use the original first column with added random noise if it is a continuous feature. For other columns, we generate $j^{th}$ column based on the j1j-1 columns of real data (treated as ${\mathbf{x}}^{ct}$) and generated data (treated as $\mathbf{x}^{\text{te}}$). We iterate this process until the last column is generated. Then for LimiX, we randomly mask a proportion of cell values of generated data and conduct missing value imputation multiple times so that we can leverage LimiX's capability of modeling the joint distribution. We conduct experiments on the following datasets: Early Stage Diabetes 2020 (Islam et al., 2019; 2020), Vertebral Column (Barreto & Neto, 2005), Seeds (Gomes Mantovani, 2015), Wine (Hardik, 2021), and Grub Damage (Vanschoren, 2014b). For evaluation, we use Trend and Shape to evaluate fidelity, which are proposed by SDMetrics [^3] https://docs.sdv.dev/sdmetrics. We also calculate the AUC of XGBoost on a hold-out test dataset using generated data or real data. From Table 25, we find that LimiX outperforms TabPFN-v2 in terms of all metrics on the five classification tasks. On the dataset of Grub Damage, the prediction performance using data generated by LimiX is even higher than using real data. The results show that LimiX is capable of capturing dependencies between features of $X$, which is brought by the pretraining strategy of mask prediction. In contrast, TabPFN-v2 is only capable of modeling $P(Y|X)$.
Table 25: Evaluation of data generation on five classification tasks. "Trend" and "Shape" are two fidelity metrics. "AUC" measures the classification performance of XGBoost using generated (or real) data. LimiX consistently outperforms TabPFN-v2 in all three metrics.
Metric Method Early Diabetes Vertebra Seeds Wine Grub Damage
Trend (↑) TabPFN-v2 0.797 0.580 0.696 0.686 0.486
LimiX 0.804 0.591 0.699 0.688 0.673
Shape (↑) TabPFN-v2 0.889 0.754 0.739 0.622 0.635
LimiX 0.902 0.763 0.768 0.646 0.762
AUC (↑) TabPFN-v2 0.839 0.652 0.861 0.670 0.695
LimiX 0.879 0.783 0.932 0.912 0.727
Real 0.969 0.915 0.982 0.996 0.710
7.8 Out-of-Distribution Generalization
In real-world applications, tabular data is often subject to distribution shifts between training data and the test data encountered during deployment. For instance, a machine learning model may be trained on patient records collected from one hospital but deployed on data from a different hospital. Such distribution shifts arising from domain variation can lead to substantial degradation in model performance. This challenge is commonly referred to as the Out-of-Distribution (OOD) generalization problem, and has been the focus of extensive research in the machine learning community (Liu et al., 2021; Yu et al., 2024).
We evaluate the OOD generalization performance of various tabular models on 10 public datasets drawn from the TableShift (Gardner et al., 2023), which is a benchmark for distribution shifts in tabular data. It contains 15 binary classification tasks in total, covering a wide range of data sources, including finance, medical diagnosis, policy, etc. More details can be found in Section A.1. To ensure fair comparisons, for each dataset, if the number of training or test samples exceeds 10,000, we randomly subsample 10,000 instances. Each experiment is repeated five times and we report the average.
From LABEL:table:tableshift_results, we can see that LimiX achieves state-of-the-art performance, securing top ranks in both ID (In-Distribution) and OOD evaluations. This could be attributed to LimiX's integration of causal data with a Context-Conditional Masked Modeling framework. This strategy enables the model to capture robust causal relationships rather than superficial correlations, leading to significantly enhanced generalization on OOD data. As for baselines, models that are also based on in-context learning (ICL), notably TabICL and TabPFN-v2, also deliver competitive results. This indicates that the ICL mechanism could better capture latent invariant patterns in data, thereby endowing models with strong generalization potential.
Table 26: Average AUC and ranks on TableShift. We can see that LimiX consistently outperforms all baselines in terms of ID or OOD performance.
Model ID_AUC ($\uparrow$) OOD_AUC ($\uparrow$) ID_Rank ($\downarrow$) OOD_Rank ($\downarrow$)
LimiX 0.848 0.806 2.5 1.3
TabICL 0.847 0.799 4.1 3.9
AutoGluon 0.842 0.797 3.5 4.0
TabPFN-v2 0.841 0.797 6.4 5.2
CatBoost 0.840 0.793 4.1 5.9
LightGBM 0.836 0.790 6.1 7.3
MLP 0.839 0.792 7.5 7.6
FT-Transformer 0.840 0.789 7.8 8.3
ResNet 0.837 0.789 8.6 8.3
NODE 0.836 0.789 9.2 9.2
XGBoost 0.830 0.783 9.9 9.5
TabDPT 0.822 0.763 12.6 12.7
TabR 0.820 0.767 13.2 13.0
TANGOS 0.801 0.752 13.8 13.6
ModernNCA 0.801 0.757 15.4 14.6
8 Conclusion
In this report, we introduce LimiX, an LDM that treats structured-data inputs as a joint distribution over variables and missingness, so that classification, regression, missing value imputation, data generation, and sample selection for interpretability, can all be expressed as conditional queries to a single model. Methodologically, LimiX adopts a lightweight, scalable architecture that models causal relations among variables while jointly capturing dependencies across features and samples. Meanwhile, LimiX combines masked joint-distribution pretraining with an episodic, context-conditional objective of per-dataset adaptation for the versatility in downstream tasks. The pretraining data for LimiX is generated with hierarchical SCMs via graph-aware and solvability-aware sampling. Attention-guided retrieval of LimiX further supports efficient inference-time ensemble and fine-tuning if desired. Empirically, experiments across 10 large structured-data benchmarks, spanning wide ranges of sample sizes, feature dimensions, number of classes, categorical-to-numerical feature ratios, missingness, and sample-to-feature ratios, confirm the effectiveness of LimiX. As a single model, LimiX consistently surpasses competitive baselines on various downstream tasks.
9 Contribution
Project Design and Lead
Xingxuan Zhang, Peng Cui
Core Contributors
Gang Ren, Han Yu, Hao Yuan, Hui Wang, Jiansheng Li, Jiayun Wu, Lang Mo, Li Mao, Mingchao Hao, Ningbo Dai, Renzhe Xu, Shuyang Li, Tianyang Zhang, Yue He, Yuanrui Wang, Yunjia Zhang, Zijing Xu
Contributors
Dongzhe Li, Fang Gao, Hao Zou, Jiandong Liu, Jiashuo Liu, Jiawei Xu, Kaijie Cheng, Kehan Li, Linjun Zhou, Qing Li, Shaohua Fan, Xiaoyu Lin, Xinyan Han, Xuanyue Li, Yan Lu, Yuan Xue, Yuanyuan Jiang, Zimu Wang, Zhenlei Wang
Appendix A Experimental Details
A.1 Details of Datasets with Distribution Shifts
The 10 public datasets we adopt from TableShift include Voting [^4] https://electionstudies.org/, Unemployment [^5] https://www.census.gov/programs-surveys/acs, Sepsis (Reyna et al., 2020), Public Health Insurance5, Income5, Hospital Readmission (Strack et al., 2014) [^6] https://archive.ics.uci.edu/ml/datasets/Diabetes+130-US+hospitals+for+years+1999-2008, Food Stamps5, Diabetes [^7] https://www.cdc.gov/brfss/index.html, College Scorecard [^8] https://collegescorecard.ed.gov/, and ASSISTments (Metz, 2020) [^9] https://new.assistments.org. Detailed information of the benchmark TableShift we adopt in Section 7.8 can be found in Table 27.
Table 27: Statistics of datasets in TableShift. "#Features" indicates the number of features. "#Train" indicates the number of training samples. "#ID_test" indicates the number of ID (In-Distribution) test samples. "#OOD_test" indicates the number of OOD (Out-of-Distribution) test samples. "Shift Type" indicates how the datasets are split into multiple domains to create distribution shifts. "#Domains" indicates the numbers of training and test domains.
Dataset #Features #Train #ID_test #OOD_test Shift type #Domains
Voting 365 34,796 4,350 21,231 Geographic Region 4/1
ASSISTments 26 2,132,526 266,566 1,906 School 386/10
Diabetes 142 969,229 121,154 209,375 Race 1/5
Food Stamps 239 629,018 78,628 48,878 Geographic Region 9/1
Hospital Readmission 183 34,288 4,287 50,968 Admission Source 15/1
Income 232 1,264,123 158,016 75,911 Geographic Region 9/1
Health Insurance 135 4,006,249 500,782 817,877 Disability Status 1/1
Sepsis 41 1,122,299 140,288 134,402 Length of Stay 47/2
Unemployment 223 1,290,914 161,365 163,611 Education Level 9/15
College ScoreCard 118 98,556 12,320 1,352 Institution Type 26/8
A.2 Hyperparameter Search Space for Baselines
The hyperparameter search space for the baseline models is summarized in Table 28.
Table 28: The hyperparameter search space for tree-based models is utilized by Optuna, which provides automatic suggestions via Bayesian optimization. The only difference between classification and regression configurations lies in the optimization target. All other hyperparameter settings remain identical.
Baseline Hyperparameter Data Type Log Search Space
RF n_estimators int no [100, 500]
max_depth int no [3, 20]
min_samples_split int no [2, 20]
min_samples_leaf int no [1, 20]
max_features categorical no {Sqrt, Log2, None}
bootstrap categorical no {True, False}
ET n_estimators int no [100, 500]
max_depth int no [3, 20]
min_samples_split int no [2, 20]
min_samples_leaf int no [1, 20]
XGBoost n_estimators int no [100, 300]
max_depth int no [3, 9]
learning_rate float yes [0.01, 0.3]
subsample float no [0.5, 1.0]
colsample_bytree float no [0.5, 1.0]
LightGBM n_estimators int no [100, 300]
learning_rate float yes [0.01, 0.3]
max_depth int no [-1, 20]
subsample float no [0.5, 1.0]
colsample_bytree float no [0.5, 1.0]
CatBoost iterations int no [100, 300]
depth int no [4, 10]
learning_rate float yes [0.01, 0.3]
Appendix B Omitted Details in Section 6
B.1 Omitted Details in Section 6.1
Proposition B.1 (Formal Version of Proposition 6.1).
Suppose that the distribution $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ has a strictly positive probability density function. Then there is a one-to-one correspondence between the distribution $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ and the family of conditionals $\{p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}):\forall\pi\in\Pi_{k}\}$.
Proof.
It is clear that $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ directly yields all conditionals $\{p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}):\pi\in\Pi_{k}\}$. We now show the converse: $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ can be recovered from this family of conditionals.
As the first step, we derive $\{p(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-j},\mathbf{X}^{\text{ct}}):j\in[d]\}$. For any $j\in[d]$, select a mask $\pi\in\Pi_{k}$ with $j\in\pi$. Then
$p(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-j},\mathbf{X}^{\text{ct}})=\frac{p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{p(\mathbf{X}^{\text{te}}_{\pi\setminus\{j\}}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}=\frac{p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{\int_{\Omega}p\left(\mathbf{X}^{\text{te}}_{\pi\setminus\{j\}},{X^{\text{te}}_{j}}^{\prime}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}\right)\mathrm{d}{X^{\text{te}}_{j}}^{\prime}}.$
Each term on the right-hand side belongs to $\{p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}):\pi\in\Pi_{k}\}$, hence $p(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-j},\mathbf{X}^{\text{ct}})$ can indeed be recovered.
We now use induction to show that the knowledge of $\{p(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-j},\mathbf{X}^{\text{ct}}):j\in[d]\}$ suffices to recover all conditionals in
${\mathcal{P}}_{k}=\{p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}):|\pi|=k\}.$
Clearly, ${\mathcal{P}}_{1}$ coincides with $\{p(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-j},\mathbf{X}^{\text{ct}}):j\in[d]\}$. Suppose ${\mathcal{P}}_{k^{\prime}}$ is obtainable for all $k^{\prime}\leq k$. Consider $p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\in{\mathcal{P}}_{k+1}$ with $|\pi|=k+1$. Pick $j\in\pi$ and set $\pi^{\prime}=\pi\setminus\{j\}$. Then
$\int_{\Omega}\frac{p(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{\pi^{\prime}},\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{p(\mathbf{X}^{\text{te}}_{\pi^{\prime}}|X^{\text{te}}_{j},\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}\,\mathrm{d}X^{\text{te}}_{j}=\int_{\Omega}\frac{p(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{p(\mathbf{X}^{\text{te}}_{\pi^{\prime}}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}\,\mathrm{d}X^{\text{te}}_{j}=\frac{1}{p(\mathbf{X}^{\text{te}}_{\pi^{\prime}}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}.$
Each term on the left-hand side belongs to ${\mathcal{P}}_{1}$ or ${\mathcal{P}}_{k}$, so $p(\mathbf{X}^{\text{te}}_{\pi^{\prime}}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})$ is obtainable. Finally,
$p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})=p(\mathbf{X}^{\text{te}}_{\pi^{\prime}}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\cdot p(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{\pi^{\prime}},\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}),$
showing that $p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})$ can be obtained as well. Hence, any element in ${\mathcal{P}}_{k+1}$ can be obtained. By induction, the claim follows. ∎
Now we show that the knowledge of a single conditional distribution $p(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-j},\mathbf{X}^{\text{ct}})$ for some $j\in[d]$ (i.e., when $\Pi=\{j\}$) is insufficient to recover the full conditional distribution $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$.
Example B.1.
Consider the case $d=2$ and $m=0$, i.e., a setting with no in-context samples and a feature dimension of $2$. In this case, the target distribution $p(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ reduces to $p(\mathbf{X}^{\text{te}})$. Suppose further that $\mathbf{X}^{\text{te}}\sim{\mathcal{N}}(\bm{\mu},\Sigma)$, where $\bm{\mu}=(\mu_{1},\mu_{2})\in{\mathbb{R}}^{2}$ and $\Sigma=\begin{pmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{pmatrix}\succeq 0$. Then the conditional distribution of $X^{\text{te}}_{1}$ given $X^{\text{te}}_{2}$ is
$X^{\text{te}}_{1}\mid X^{\text{te}}_{2}\sim{\mathcal{N}}\left(\Sigma_{12}\Sigma_{22}^{-1}X^{\text{te}}_{2}+\mu_{1}-\Sigma_{12}\Sigma_{22}^{-1}\mu_{2},\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{-1}\Sigma_{21}\right).$
Consequently, knowledge of $p(X^{\text{te}}_{1}|X^{\text{te}}_{2})$ alone provides access only to the quantities $\Sigma_{12}\Sigma_{22}^{-1}$, $\mu_{1}-\Sigma_{12}\Sigma_{22}^{-1}\mu_{2}$, and $\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{-1}\Sigma_{21}$. These are insufficient to uniquely determine the full parameter set $(\bm{\mu},\Sigma)$, and therefore $p(\mathbf{X}^{\text{te}})$ cannot be fully recovered from $p(X^{\text{te}}_{1}|X^{\text{te}}_{2})$ alone. By symmetry, $p(\mathbf{X}^{\text{te}})$ also cannot be recovered solely from $p(X^{\text{te}}_{2}|X^{\text{te}}_{1})$.
B.2 Omitted Details in Section 6.2
We denote by $q_{\theta}(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}}),\theta\in\Theta$ the distribution induced by the learned family of conditional probabilities $\{q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}):\pi\in\Pi_{k}\}$. We overload the notation by writing $q_{\theta}(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\coloneqq q_{\theta}(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})p(\mathbf{X}^{\text{ct}})$.
B.2.1 Sample Efficiency
Theorem B.2 (Formal Version of Theorem 6.2).
Suppose there exists $\theta^{*}\in\Theta$ such that $q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})=p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})$ for all $\mathbf{X}^{\text{ct}}\in\Omega^{m\times d}$, $\mathbf{X}^{\text{te}}\in\Omega^{d}$, and $\pi\subseteq[d]$, and that the minimizer of $L_{k}(\theta)$ is unique for every $k$. Assume that for all $\theta\in\Theta$, $\mathbf{X}^{\text{ct}}\in\Omega^{m\times d}$, $\mathbf{X}^{\text{te}}\in\Omega^{d}$, and $\pi\subseteq[d]$, the gradient norm $\lVert\nabla_{\theta}q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\rVert_{2}$ and the Hessian norm $\lVert\nabla^{2}_{\theta}q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\rVert_{F}$ exist and are finite. Assume $\nabla^{2}_{\theta}L_{k}(\theta^{*})\succ 0$ for all $k\in[d]$. Then, for every sufficiently small neighborhood $\mathcal{B}$} of $\theta^{*}$, there exists a sufficiently large $n$ such that $\hat{\theta}_{k,n}$ is the unique minimizer of $\hat{L}_{k}(\theta)$ in $\mathcal{B}$}. Moreover,
$\sqrt{n}(\hat{\theta}_{k,n}-\theta^{*})\xrightarrow{d}{\mathcal{N}}(0,\Gamma_{k}),$
where $\Gamma_{k}$ does not depend on $n$ and satisfies $\Gamma_{k+1}\preceq\Gamma_{k}$.
The proof of Theorem B.2 is based on the following lemma.
Lemma B.3 (Van der Vaart (2000), Theorem 5.23; statement adapted from Qin & Risteski (2024); Li et al. (2024c)).
Consider a loss $L:\Theta\rightarrow\mathbb{R}$, such that $L(\theta)=\mathbb{E}_{\mathbf{X}\sim p}[\ell_{\theta}(\mathbf{X})]$ for $\ell_{\theta}:{\mathcal{X}}\rightarrow{\mathbb{R}}$. Let $\Theta^{*}$ be the the set of global minima of $L$, i.e.,
$\Theta^{*}=\{\theta^{*}:L(\theta^{*})=\min_{\theta\in\Theta}L(\theta)\}.$
Suppose the following conditions are met:
  • (Gradient bounds on $\ell_{\theta}$) The map $\theta\mapsto\ell_{\theta}$ is measurable and differentiable at every $\theta^{*}\in\Theta^{*}$ for $p$-almost every $\mathbf{X}$. Furthermore, there exists a function $B(\mathbf{X})$, s.t. $\mathbb{E}[B(\mathbf{X})^{2}]<\infty$ and and for every $\theta_{1},\theta_{2}$ near $\theta^{*}$, we have
    $\left|\ell_{\theta_{1}}(\mathbf{X})-\ell_{\theta_{2}}(\mathbf{X})\right|<B(\mathbf{X})\left\lVert\theta_{1}-\theta_{2}\right\rVert_{2}$
  • (Twice-differentiability of $L$) $L(\theta)$ is twice-differentiable at every $\theta^{*}\in\Theta^{*}$ with Hessian $\nabla^{2}_{\theta}L(\theta^{*})$, and furthermore $\nabla^{2}_{\theta}L(\theta^{*})\succ 0$.
  • (Uniform law of large numbers) The loss $L$ satisfies a uniform law of large numbers, that is
    $\sup_{\theta\in\Theta}\left|\hat{\mathbb{E}}[\ell_{\theta}(\mathbf{X})]-L(\theta)\right|\xrightarrow{p}0.$
  • (Realizability) The data distribution $p$ satisfies: $\exists\theta^{*}\in\Theta$ such that $p_{\theta^{*}}=p$.
Then for every $\theta^{*}\in\Theta^{*}$, and every sufficiently small neighborhood $S$ of $\theta^{*}$, there exists a sufficiently large $n$, such that there is a unique minimizer $\hat{\theta}_{n}$ of $\hat{\mathbb{E}}[\ell_{\theta}(\mathbf{X})]$ in $S$. Furthermore, $\hat{\theta}_{n}$ satisfies:
$\sqrt{n}\left(\hat{\theta}_{n}-\theta^{*}\right)\xrightarrow{d}{\mathcal{N}}\left(0,\left(\nabla_{\theta}^{2}L\left(\theta^{*}\right)\right)^{-1}\operatorname{Cov}\left(\nabla_{\theta}\ell_{\theta^{*}}\left(\mathbf{X}\right)\right)\left(\nabla_{\theta}^{2}L\left(\theta^{*}\right)\right)^{-1}\right).$
Similar to Lemma 2 in Li et al. (2024c), we have the following lemma.
Lemma B.4.
Under the same assumptions as in Theorem B.2, we have
$\nabla_{\theta}^{2}L_{k}(\theta^{*})=\mathrm{Cov}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\,\pi\sim\mathrm{Unif}(\Pi_{k})}\left(-\nabla_{\theta}\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\pi}\mid\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}\right)\right).$ (5)
Proof.
It is easy to verify that $\theta^{*}$ is the minimizer of $L_{k}(\theta)$ for any $k\in[d]$ and hence $\theta^{*}_{k}=\theta^{*}$ by assumption.
Rewrite $\nabla_{\theta}^{2}L_{k}(\theta^{*})$ and we get that
$\,\nabla_{\theta}^{2}L_{k}(\theta^{*})$ (6)
$=$ $\,\nabla_{\theta}^{2}\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\pi}\left[-\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right]$
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\pi}\left[-\nabla_{\theta}^{2}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right]$
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\pi}\left[\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)^{\top}-\frac{\nabla_{\theta}^{2}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}\right].$
In addition, we have
$\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\pi}\left[\frac{\nabla_{\theta}^{2}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}\right]=\mathbb{E}_{\mathbf{X}^{\text{ct}},\pi}\mathbb{E}_{\mathbf{X}^{\text{te}}_{-\pi}|\mathbf{X}^{\text{ct}}}\mathbb{E}_{\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}}\left[\frac{\nabla_{\theta}^{2}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}\right]$
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\pi}\mathbb{E}_{\mathbf{X}^{\text{te}}_{-\pi}|\mathbf{X}^{\text{ct}}}\left[\int_{\Omega^{k}}\frac{\nabla_{\theta}^{2}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}\cdot p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\mathrm{d}\mathbf{X}^{\text{te}}_{\pi}\right]$
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\pi}\mathbb{E}_{\mathbf{X}^{\text{te}}_{-\pi}|\mathbf{X}^{\text{ct}}}\left[\int_{\Omega^{k}}\nabla_{\theta}^{2}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\mathrm{d}\mathbf{X}^{\text{te}}_{\pi}\right]$ (Due to the assumption $q_{\theta^{*}}=p$)
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\pi}\mathbb{E}_{\mathbf{X}^{\text{te}}_{-\pi}|\mathbf{X}^{\text{ct}}}\left[\nabla_{\theta}^{2}\int_{\Omega^{k}}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\mathrm{d}\mathbf{X}^{\text{te}}_{\pi}\right]=\mathbb{E}_{\mathbf{X}^{\text{ct}},\pi}\mathbb{E}_{\mathbf{X}^{\text{te}}_{-\pi}|\mathbf{X}^{\text{ct}}}\left[\nabla_{\theta}^{2}1\right]=0.$
Combined with Equation 6, we can get that
$\nabla_{\theta}^{2}L_{k}(\theta^{*})=\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\pi}\left[\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)^{\top}\right].$ (7)
Now rewrite the right-hand side of Equation 5 and we get
$\begin{aligned} &\,\mathrm{Cov}\left(-\nabla_{\theta}\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\pi}\mid\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}\right)\right)\\ =&\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\pi}\left[\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)^{\top}\right]\\ &\,-\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\pi}\left[\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right]\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\pi}\left[\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right]^{\top}\end{aligned}.$ (8)
Note that
$\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\pi}\left[\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right]=\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\pi}\left[\frac{\nabla_{\theta}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}\right]$
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\pi}\mathbb{E}_{\mathbf{X}^{\text{te}}_{-\pi}|\mathbf{X}^{\text{ct}}}\mathbb{E}_{\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}}\left[\frac{\nabla_{\theta}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}\right]$
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\pi}\mathbb{E}_{\mathbf{X}^{\text{te}}_{-\pi}|\mathbf{X}^{\text{ct}}}\left[\int_{\Omega^{k}}\frac{\nabla_{\theta}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}{q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})}\cdot p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\mathrm{d}\mathbf{X}^{\text{te}}_{\pi}\right]$
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\pi}\mathbb{E}_{\mathbf{X}^{\text{te}}_{-\pi}|\mathbf{X}^{\text{ct}}}\left[\int_{\Omega^{k}}\nabla_{\theta}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\mathrm{d}\mathbf{X}^{\text{te}}_{\pi}\right]$ (Due to the assumption $q_{\theta^{*}}=p$)
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\pi}\mathbb{E}_{\mathbf{X}^{\text{te}}_{-\pi}|\mathbf{X}^{\text{ct}}}\left[\nabla_{\theta}\int_{\Omega^{k}}q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\mathrm{d}\mathbf{X}^{\text{te}}_{\pi}\right]$
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\pi}\mathbb{E}_{\mathbf{X}^{\text{te}}_{-\pi}|\mathbf{X}^{\text{ct}}}\left[\nabla_{\theta}1\right]=0.$
Combined with Equation 8, we can get that
$\mathrm{Cov}\left(-\nabla_{\theta}\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\pi}\mid\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}\right)\right)=\mathbb{E}\left[\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)^{\top}\right].$ (9)
Now the claim follows from Equations 7 and 9. ∎
Based on the above lemmas, we can now prove Theorem B.2. The proof follows a similar idea to that of Theorem 1 in Li et al. (2024c).
Proof of Theorem B.2.
It is easy to verify that $\theta^{*}$ is the minimizer of $L_{k}(\theta)$ for any $k\in[d]$ and hence $\theta^{*}_{k}=\theta^{*}$ by assumption. According to Lemma B.3, it holds that for every sufficiently small neighborhood $S$ of $\theta^{*}$, there exists a sufficiently large $n$, such that there is a unique minimizer $\hat{\theta}_{k,n}$ in $S$. Furthermore, $\hat{\theta}_{k,n}$ satisfies:
$\sqrt{n}\left(\hat{\theta}_{k,n}-\theta^{*}\right)\xrightarrow{d}{\mathcal{N}}\left(0,\Gamma_{k}\right).$
Here due to Lemma B.4, it holds that
$\Gamma_{k}$ $=\left(\nabla_{\theta}^{2}L_{k}\left(\theta^{*}\right)\right)^{-1}\mathrm{Cov}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\,\pi\sim\mathrm{Unif}(\Pi_{k})}\left(-\nabla_{\theta}\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}}\right)\right)\left(\nabla_{\theta}^{2}L_{k}\left(\theta^{*}\right)\right)^{-1}$ (10)
$=\left(\nabla_{\theta}^{2}L_{k}\left(\theta^{*}\right)\right)^{-1}.$
Fix a $k\in[d-1]$. Then for every $\pi\in\Pi_{k+1}$ and $j\in\pi$, let $\gamma=\pi\setminus\{j\}$ and we have
$\log q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})$ $=\log q_{\theta}\left(\mathbf{X}^{\text{te}}_{\gamma},X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-(\gamma\cup\{j\})},\mathbf{X}^{\text{ct}}\right)$ (11)
$=\log q_{\theta}\left(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-(\gamma\cup\{j\})},\mathbf{X}^{\text{ct}}\right)+\log q_{\theta}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right).$
As a result, we have
$\,\nabla^{2}_{\theta}L_{k+1}(\theta^{*})$
$=$ $\,\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\pi\sim\mathrm{Unif}(\Pi_{k+1})}\left[\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)^{\top}\right]$ (Due to Equation 7)
$=$ $\,\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\gamma\sim\mathrm{Unif}(\Pi_{k}),j\in\mathrm{Unif}([d]\setminus\gamma)}\left[\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)\left(\nabla_{\theta}\log q_{\theta^{*}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{te}}_{-\pi},\mathbf{X}^{\text{ct}})\right)^{\top}\right]$ (Letting $\pi=\gamma\cup\{j\}$)
$=$ $\,\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\gamma,j}\Bigg{[}\left(\nabla_{\theta}\log q_{\theta^{*}}\left(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-(\gamma\cup\{j\})},\mathbf{X}^{\text{ct}}\right)+\nabla_{\theta}\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\right)\times$
$\qquad\left(\nabla_{\theta}\log q_{\theta^{*}}\left(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-(\gamma\cup\{j\})},\mathbf{X}^{\text{ct}}\right)+\nabla_{\theta}\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\right)^{\top}\Bigg{]}.$ (By Equation 11)
Define
$A$ $=\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\gamma,j}\left[\left(\nabla_{\theta}\log q_{\theta^{*}}\left(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-(\gamma\cup\{j\})},\mathbf{X}^{\text{ct}}\right)\right)\left(\nabla_{\theta}\log q_{\theta^{*}}\left(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-(\gamma\cup\{j\})},\mathbf{X}^{\text{ct}}\right)\right)^{\top}\right].$
$B$ $=\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\gamma,j}\left[\left(\nabla_{\theta}\log q_{\theta^{*}}\left(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-(\gamma\cup\{j\})},\mathbf{X}^{\text{ct}}\right)\right)\left(\nabla_{\theta}\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\right)^{\top}\right].$
$C$ $=\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\gamma,j}\left[\left(\nabla_{\theta}\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\right)\left(\nabla_{\theta}\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\right)^{\top}\right].$
Then we have
$\nabla^{2}_{\theta}L_{k+1}(\theta^{*})=A+B+B^{\top}+C.$
It is easy to verify that $C=\nabla^{2}_{\theta}L_{k}(\theta^{*})$. Now consider $B$. It holds that
$B$ $=\mathbb{E}_{\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}},\gamma,j}\left[\left(\nabla\log q_{\theta^{*}}\left(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-(\gamma\cup\{j\})},\mathbf{X}^{\text{ct}}\right)\right)\left(\nabla\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\right)^{\top}\right]$
$=\mathbb{E}_{\mathbf{X}^{\text{ct}},\gamma,j,\mathbf{X}^{\text{te}}_{-\gamma}}\left[\left(\nabla\log q_{\theta^{*}}\left(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-(\gamma\cup\{j\})},\mathbf{X}^{\text{ct}}\right)\right)\cdot\mathbb{E}_{\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}}\left[\left(\nabla\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\right)^{\top}\right]\right]$
$=\mathbb{E}_{\mathbf{X}^{\text{ct}},\gamma,j,\mathbf{X}^{\text{te}}_{-\gamma}}\left[\left(\nabla\log q_{\theta^{*}}\left(X^{\text{te}}_{j}|\mathbf{X}^{\text{te}}_{-(\gamma\cup\{j\})},\mathbf{X}^{\text{ct}}\right)\right)\cdot 0\right]$ (See Equation 12 below)
$=0.$
Here the third equation is due to:
$\mathbb{E}_{\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}}\left[\nabla_{\theta}\log q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\right]$ $=\int_{\Omega^{k}}\frac{\nabla_{\theta}q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)}{q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)}\cdot p\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\mathrm{d}\mathbf{X}^{\text{te}}_{\gamma}$ (12)
$=\int_{\Omega^{k}}\nabla_{\theta}q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\mathrm{d}\mathbf{X}^{\text{te}}_{\gamma}$
$=\nabla_{\theta}\int_{\Omega^{k}}q_{\theta^{*}}\left(\mathbf{X}^{\text{te}}_{\gamma}|\mathbf{X}^{\text{te}}_{-\gamma},\mathbf{X}^{\text{ct}}\right)\mathrm{d}\mathbf{X}^{\text{te}}_{\gamma}=\nabla_{\theta}1=0.$
In addition, since $A\succeq 0$, we have
$\nabla^{2}_{\theta}L_{k+1}(\theta^{*})=A+B+B^{\top}+C=A+\nabla^{2}_{\theta}L_{k}(\theta^{*})\succeq\nabla^{2}_{\theta}L_{k}(\theta^{*}).$
Noting that $\Gamma_{k}=\big{(}\nabla^{2}_{\theta}L_{k}(\theta^{*})\big{)}^{-1}$ by Equation 10, it follows that $\Gamma_{k+1}\preceq\Gamma_{k}$. Now the claim follows. ∎
B.2.2 Generalization for Joint Distribution Learning
The generalization error for joint distribution learning is characterized by a key concept termed approximate tensorization of entropy. It measures the ``complexity'' of the distribution over $(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ by evaluating how easily an algorithm can generate samples from the joint distribution $q(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ with access to local conditional distributions $q(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})$. Technically, approximate tensorization of entropy is associated with the mixing time of Gibbs sampling dynamics, which is the sample generation algorithm to be considered.
Definition B.1 (Approximate Tensorization of Entropy (Caputo & Parisi, 2021)).
For a distribution $q(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ and the set of $k$-cell masks $\Pi_{k}$, if there exists a constant $C_{k}(q)$ depending on $q$ and $k$, such that for any distribution $r$ over $(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$,
$D_{\operatorname{KL}}(r\parallel q)\leq C_{k}(q)\cdot\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\pi\sim\mathrm{Unif}(\Pi_{k})}\left[D_{\operatorname{KL}}(r(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})\parallel q(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi}))\right],$
then $q$ satisfies approximate tensorization of entropy with respect to the constant $C_{k}$ and the mask set $\Pi_{k}$. Let $\underline{C_{k}}(q)$ be the minimum of all possible constants $C_{k}(q)$ such that $q$ satisfies approximate tensorization of entropy.
Before presenting the main result, we introduce a few regularity conditions in the parametric class $q_{\theta}(\mathbf{X}^{\text{te}}|\mathbf{X}^{\text{ct}})$ that defines the model.
Assumption B.1 (Regularity Conditions in the Parametric Class (Li et al., 2024c)).
  1. 1.
    There exists $\beta\in(0,1)$ such that $\forall\leavevmode\nobreak\ 1\leq k\leq d,\forall\leavevmode\nobreak\ \pi\in\Pi_{k}$ and $\forall\leavevmode\nobreak\ \theta\in\Theta$, $p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})>0$ implies $q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})>\beta$.
  2. 2.
    For any $\epsilon>0$, there exists a partition $\text{Par}(\Theta)=\{\Theta_{1},\dots,\Theta_{|\text{Par}(\Theta)|}\}$ of $\Theta$, such that $\forall\leavevmode\nobreak\ 1\leq k\leq d,\forall\leavevmode\nobreak\ \pi\in\Pi_{k}$, $\forall\leavevmode\nobreak\ \Theta_{i}\in\text{Par}(\Theta)$, $\forall\leavevmode\nobreak\ \theta_{1},\theta_{2}\in\Theta_{i}$, and any $(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$,
    $\left|\log q_{\theta_{1}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})-\log q_{\theta_{2}}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})\right|\leq\frac{\epsilon}{2}.$
    Let $N(\Theta,\epsilon)$ be the cardinality of the smallest partition $\text{Par}(\Theta)$ that satisfies the condition above.
The first assumption implies that the true distribution $p(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})$ supports the parametric distribution $q_{\theta}(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})$ . The second assumption specifies the covering number of the parameter space $\Theta$ and the lipschitz continuity of the log-likelihood loss function.
Theorem B.5 (Formal Version of Theorem 6.3).
For $\theta\in\Theta$, assume $q_{\theta}(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ satisfies approximate tensorization of entropy with respect to some constant $C_{1}(q_{\theta})$ and the mask set $\Pi_{1}$. Then for any $1\leq k\leq d$, $q_{\theta}(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ satisfies approximate tensorization of entropy with respect to some constant $C_{k}(q_{\theta})$ and the mask set $\Pi_{k}$. Furthermore, $\underline{C_{k+1}}(q_{\theta})\leq\underline{C_{k}}(q_{\theta})$.
Under Assumption B.1, for any $\epsilon>0$ and any $\delta\in(0,\frac{1}{d})$, with probability at least $1-d\delta$, for any $1\leq k\leq d$ and any $\theta\in\Theta$,
$\mathbb{E}_{\mathbf{X}^{\text{ct}}\sim p}\left[D_{\operatorname{TV}}(q_{\theta}(\cdot|\mathbf{X}^{\text{ct}})\parallel p(\cdot|\mathbf{X}^{\text{ct}}))\right]<\sqrt{\frac{1}{2}\underline{C_{k}}(q_{\theta})\left(\hat{L}_{k}(\theta)+B\log\frac{1}{\beta}+\epsilon\right)+C},$
where $B=\sqrt{\frac{1}{\delta}\cdot(8|\Omega|)^{d(m+1)}N(\Theta,\epsilon)}+\sqrt{\frac{1}{2n}\cdot\log\frac{8N(\Theta,\epsilon)}{\delta}}$, and $C=\sqrt{\frac{|\Omega|^{3d(m+1)}}{8\delta n}}$.
With a sufficiently large sample size $n$ and a sufficiently small loss value $\hat{L}_{k}(\theta)$, the upper bound is dominated by the term $\sqrt{\frac{1}{2}\underline{C_{k}}(q_{\theta})(B\log\frac{1}{\beta}+\epsilon)}$, which is scaled by $\underline{C_{k}}(q_{\theta})$, the constant for the approximate tensorization of entropy. The theorem implies a reduced upper bound for the estimation error of the joint distribution with an increasing number of masked cells.
We prove the monotonicity of $\underline{C_{k}}(q_{\theta})$ with respect to $k$ in Proposition B.6, and prove the upper bound for generalization error in Proposition B.7.
Proposition B.6.
For $\theta\in\Theta$, assume $q_{\theta}(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ satisfies approximate tensorization of entropy with respect to some constant $C_{1}(q_{\theta})$ and the mask set $\Pi_{1}$. Then for any $1\leq k\leq d$, $q_{\theta}(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ satisfies approximate tensorization of entropy with respect to some constant $C_{k}(q_{\theta})$ and the mask set $\Pi_{k}$. Furthermore, $\underline{C_{k+1}}(q_{\theta})\leq\underline{C_{k}}(q_{\theta})$.
Proof of Proposition B.6.
Since $q_{\theta}(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ satisfies approximate tensorization of entropy with respect to some constant $C_{1}(q_{\theta})$ and the mask set $\Pi_{1}$, for any distribution $r$ over $(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$,
$D_{\operatorname{KL}}(r\parallel q_{\theta})\leq C_{1}(q_{\theta})\cdot\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\pi\sim\mathrm{Unif}(\Pi_{1})}\left[D_{\operatorname{KL}}(r(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})\parallel q_{\theta}(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi}))\right].$
By Definition B.1,
$D_{\operatorname{KL}}(r\parallel q_{\theta})\leq\underline{C_{1}}(q_{\theta})\cdot\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\pi\sim\mathrm{Unif}(\Pi_{1})}\left[D_{\operatorname{KL}}(r(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})\parallel q_{\theta}(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi}))\right].$
We prove the proposition by deduction. Assume for some $1\leq k<d$, $q_{\theta}(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ satisfies approximate tensorization of entropy with respect to some constant $C_{k}(q_{\theta})$ and the mask set $\Pi_{k}$. It follows that
$D_{\operatorname{KL}}(r\parallel q_{\theta})\leq\underline{C_{k}}(q_{\theta})\cdot\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\pi\sim\mathrm{Unif}(\Pi_{k})}\left[D_{\operatorname{KL}}(r(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})\parallel q_{\theta}(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi}))\right].$ (13)
We have
$\begin{split}&\quad\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\pi\sim\mathrm{Unif}(\Pi_{k+1})}\left[D_{\operatorname{KL}}(r(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})\parallel q_{\theta}(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi}))\right]\\ &=\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\pi\sim\mathrm{Unif}(\Pi_{k}),a\sim\mathrm{Unif}([d]\backslash\pi)}\left[D_{\operatorname{KL}}(r(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi\cup\{a\}})\parallel q_{\theta}(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi\cup\{a\}}))\right]\\ &\geq\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\pi\sim\mathrm{Unif}(\Pi_{k}),a\sim\mathrm{Unif}([d]\backslash\pi)}\left[D_{\operatorname{KL}}(r(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})\parallel q_{\theta}(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi}))\right]\\ &=\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\pi\sim\mathrm{Unif}(\Pi_{k})}\left[D_{\operatorname{KL}}(r(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})\parallel q_{\theta}(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi}))\right].\end{split}$ (14)
The inequality follows from the data processing inequality:
$D_{\operatorname{KL}}(p(x,y)\parallel q(x,y))=\mathbb{E}_{x}\left[D_{\operatorname{KL}}(p(y|x)\parallel q(y|x))\right]+D_{\operatorname{KL}}(p(x)\parallel q(x))\geq\mathbb{E}_{x}\left[D_{\operatorname{KL}}(p(y|x)\parallel q(y|x))\right].$
Combining Equations 13 and LABEL:eq:app_monotonous_2,
$D_{\operatorname{KL}}(r\parallel q_{\theta})\leq\underline{C_{k}}(q_{\theta})\cdot\mathbb{E}_{(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})\sim p,\pi\sim\mathrm{Unif}(\Pi_{k+1})}\left[D_{\operatorname{KL}}(r(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})\parallel q_{\theta}(\cdot|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi}))\right].$
Therefore, $q_{\theta}(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ satisfies approximate tensorization of entropy with respect to $\underline{C_{k}}(q_{\theta})$ and the mask set $\Pi_{k+1}$. It follows that $\underline{C_{k+1}}(q_{\theta})\leq\underline{C_{k}}(q_{\theta})$. By deduction, for any $1\leq k\leq d$, $q_{\theta}(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$ satisfies approximate tensorization of entropy with respect to some constant $C_{k}(q_{\theta})$ and the mask set $\Pi_{k}$. ∎
Proposition B.7.
Under Assumption B.1 and the condition in Proposition B.6, for any $\epsilon>0$ and any $\delta\in(0,\frac{1}{d})$, with probability at least $1-d\delta$, for any $1\leq k\leq d$ and any $\theta\in\Theta$,
$\mathbb{E}_{\mathbf{X}^{\text{ct}}\sim p}\left[D_{\operatorname{TV}}(q_{\theta}(\cdot|\mathbf{X}^{\text{ct}})\parallel p(\cdot|\mathbf{X}^{\text{ct}}))\right]<\sqrt{\frac{1}{2}\underline{C_{k}}(q_{\theta})\left(\hat{L}_{k}(\theta)+B\log\frac{1}{\beta}+\epsilon\right)+C},$
where $B=\sqrt{\frac{1}{\delta}\cdot(8|\Omega|)^{d(m+1)}N(\Theta,\epsilon)}+\sqrt{\frac{1}{2n}\cdot\log\frac{8N(\Theta,\epsilon)}{\delta}}$, and $C=\sqrt{\frac{|\Omega|^{3d(m+1)}}{8\delta n}}$.
The proposition is a corollary from Theorem 4 in Li et al. (2024c), which provides an upper bound for $D_{\operatorname{TV}}(q_{\theta}\parallel p)$ in our setting. The remaining gap is an extension of the result to total variation between conditional distributions. We present Li et al. (2024c)'s result as a lemma.
Lemma B.8 (Li et al. (2024c), Theorem 4).
Consider random variables $\mathbf{X}\in\Omega^{d}$ and $\pi\subset[d]$. We are given $n$ i.i.d. samples $\{{\mathbf{x}}^{(i)}\}_{i=1}^{n}$ drawn from $p(\mathbf{X})$. For each sample ${\mathbf{x}}^{(i)}$, we observe $l$ i.i.d. masks $\{\pi_{j}^{(i)}\}_{j=1}^{l}$ drawn from $p(\pi|{\mathbf{x}}^{(i)})$. Consider the empirical loss function:
$\hat{L}(\theta)=\frac{1}{nl}\sum_{i=1}^{n}\sum_{j=1}^{l}-\log q_{\theta}\left({\mathbf{x}}_{\pi_{j}^{(i)}}^{(i)}|{\mathbf{x}}_{-\pi_{j}^{(i)}}^{(i)},\pi_{j}^{(i)}\right).$
Suppose the following conditions are met:
  1. 1.
    There exists a constant $C(q_{\theta})$ depending on $q_{\theta}$, such that for any distribution $r$ over $\mathbf{X}$,
    $D_{\operatorname{KL}}(r\parallel q)\leq C(q_{\theta})\cdot\mathbb{E}_{\mathbf{X}\sim p(\mathbf{X}),\pi\sim p(\pi|\mathbf{X})}\left[D_{\operatorname{KL}}(r(\cdot|\mathbf{X}_{-\pi},\pi)\parallel q(\cdot|\mathbf{X}_{-\pi},\pi)\right].$
  2. 2.
    There exists $\beta\in(0,1)$ such that $\forall\leavevmode\nobreak\ \pi\subset[d]$ and $\forall\leavevmode\nobreak\ \theta\in\Theta$, $p({\mathbf{x}}_{\pi}|{\mathbf{x}}_{-\pi},\pi)>0$ implies $q_{\theta}({\mathbf{x}}_{\pi}|{\mathbf{x}}_{-\pi},\pi)>\beta$.
  3. 3.
    For any $\epsilon>0$, there exists a partition $\text{Par}(\Theta)=\{\Theta_{1},\dots,\Theta_{|\text{Par}(\Theta)|}\}$ of $\Theta$, such that $\forall\leavevmode\nobreak\ \pi\subset[d]$, $\forall\leavevmode\nobreak\ \Theta_{i}\in\text{Par}(\Theta)$, $\forall\leavevmode\nobreak\ \theta_{1},\theta_{2}\in\Theta_{i}$, and any $\mathbf{x}$},
    $\left|\log q_{\theta_{1}}({\mathbf{x}}_{\pi}|{\mathbf{x}}_{-\pi},\pi)-\log q_{\theta_{2}}({\mathbf{x}}_{\pi}|{\mathbf{x}}_{-\pi},\pi)\right|\leq\frac{\epsilon}{2}.$
    Let $N(\Theta,\epsilon)$ be the cardinality of the smallest partition $\text{Par}(\Theta)$ that satisfies the condition above.
Then for any $\epsilon>0$ and any $\delta\in(0,1)$, with probability at least $1-\delta$, for any $\theta\in\Theta$,
$D_{\operatorname{TV}}(q_{\theta}(\mathbf{X})\parallel p(\mathbf{X}))<\sqrt{\frac{1}{2}C(q_{\theta})\left(\hat{L}(\theta)+B\log\frac{1}{\beta}+\epsilon\right)+C},$
where $B=\sqrt{\frac{1}{l\delta}\cdot(8|\Omega|)^{d}N(\Theta,\epsilon)}+\sqrt{\frac{1}{2n}\cdot\log\frac{8N(\Theta,\epsilon)}{\delta}}$, and $C=\sqrt{\frac{|\Omega|^{3d}}{8\delta n}}$.
Remark B.1.
Theorem 4 in Li et al. (2024c) is specified for $\hat{\theta}$ as the minimizer of the empirical loss function. In fact, the proof applies uniformly to arbitrarily $\theta\in\Theta$.
Proof of Proposition B.7.
For each pair of $(\mathbf{x}^{\text{ct},(i)},\mathbf{x}^{\text{te},(i)})$, exactly one mask $\pi_{i}$ is drawn independently from $\mathrm{Unif}(\Pi_{k})$. Therefore, $\pi\perp\mkern-9.0mu\perp(\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}})$. It follows from Lemma B.8 that for each $1\leq k\leq d$, for any $\epsilon>0$ and any $\delta\in(0,1)$, with probability at least $1-\delta$, for any $\theta\in\Theta$,
$D_{\operatorname{TV}}(q_{\theta}(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})\parallel p(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}}))<\sqrt{\frac{1}{2}\underline{C_{k}}(q_{\theta})\left(\hat{L}_{k}(\theta)+B\log\frac{1}{\beta}+\epsilon\right)+C},$ (15)
where $B=\sqrt{\frac{1}{\delta}\cdot(8|\Omega|)^{d(m+1)}N(\Theta,\epsilon)}+\sqrt{\frac{1}{2n}\cdot\log\frac{8N(\Theta,\epsilon)}{\delta}}$, and $C=\sqrt{\frac{|\Omega|^{3d(m+1)}}{8\delta n}}$.
By union bound, for any $\epsilon>0$ and any $\delta\in(0,\frac{1}{d})$, with probability at least $1-d\delta$, Equation 15 is satisfied for any $1\leq k\leq d$ and any $\theta\in\Theta$.
Furthermore, we have
$\begin{split}\mathbb{E}_{\mathbf{X}^{\text{ct}}\sim p}\left[D_{\operatorname{TV}}(q_{\theta}(\cdot|\mathbf{X}^{\text{ct}})\parallel p(\cdot|\mathbf{X}^{\text{ct}}))\right]&=\sum_{\mathbf{x}^{\text{ct}}\in\Omega^{m\times d}}p(\mathbf{x}^{\text{ct}})D_{\operatorname{TV}}(q_{\theta}(\cdot|\mathbf{X}^{\text{ct}}=\mathbf{x}^{\text{ct}})\parallel p(\cdot|\mathbf{X}^{\text{ct}}=\mathbf{x}^{\text{ct}}))\\ &=\sum_{\mathbf{x}^{\text{ct}}\in\Omega^{m\times d}}p(\mathbf{x}^{\text{ct}})\cdot\frac{1}{2}\sum_{\mathbf{x}^{\text{te}}\in\Omega^{d}}\left|q_{\theta}(\mathbf{x}^{\text{te}}|\mathbf{x}^{\text{ct}})-p(\mathbf{x}^{\text{te}}|\mathbf{x}^{\text{ct}})\right|\\ &=\frac{1}{2}\sum_{\mathbf{x}^{\text{ct}}\in\Omega^{m\times d},\mathbf{x}^{\text{te}}\in\Omega^{d}}\left|p(\mathbf{x}^{\text{ct}})q_{\theta}(\mathbf{x}^{\text{te}}|\mathbf{x}^{\text{ct}})-p(\mathbf{x}^{\text{ct}})p(\mathbf{x}^{\text{te}}|\mathbf{x}^{\text{ct}})\right|\\ &=\frac{1}{2}\sum_{\mathbf{x}^{\text{ct}}\in\Omega^{m\times d},\mathbf{x}^{\text{te}}\in\Omega^{d}}\left|q_{\theta}(\mathbf{x}^{\text{te}},\mathbf{x}^{\text{ct}})-p(\mathbf{x}^{\text{te}},\mathbf{x}^{\text{ct}})\right|\\ &=D_{\operatorname{TV}}(q_{\theta}(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})\parallel p(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})).\end{split}$ (16)
The proof is complete by combining Equations 15 and 16. ∎